### v1

# SurgiATM: A Physics-Guided Plug-and-Play Model for Deep Learning-Based Smoke Removal in Laparoscopic Surgery

Abstract
During laparoscopic surgery, smoke generated by tissue cauterization can significantly degrade the visual quality of endoscopic frames, increasing the risk of surgical errors and hindering both clinical decision-making and computer-assisted visual analysis. Consequently, removing surgical smoke is critical to ensuring patient safety and maintaining operative efficiency. In this study, we propose the Surgical Atmospheric Model (SurgiATM) for surgical smoke removal. SurgiATM statistically bridges a physics-based atmospheric model and data-driven deep learning models, combining the superior generalizability of the former with the high accuracy of the latter. Furthermore, SurgiATM is designed as a lightweight, plug-and-play module that can be seamlessly integrated into diverse surgical desmoking architectures to enhance their accuracy and stability, better meeting clinical requirements. It introduces only two hyperparameters and no additional trainable weights, preserving the original network architecture with minimal computational and modification overhead. We conduct extensive experiments on three public surgical datasets with ten desmoking methods, involving multiple network architectures and covering diverse procedures, including cholecystectomy, partial nephrectomy, and diaphragm dissection. The results demonstrate that incorporating SurgiATM commonly reduces the restoration errors of existing models and relatively enhances their generalizability, without adding any trainable layers or weights. This highlights the convenience, low cost, effectiveness, and generalizability of the proposed method. The code for SurgiATM is released at https://github.com/MingyuShengSMY/SurgiATM.
{IEEEkeywords}
Endoscopy, laparoscopic surgery, smoke removal, plug-and-play, parameter-efficient
1 Introduction
[…]
Minimally invasive surgery (MIS) has noticeably benefited patients by reducing risk and enabling faster recovery [Maier_Hein_2017_Surgical_data, Maier_Hein_2022_Surgical_data]. However, surgical smoke, commonly generated by high-energy instruments (e.g., electrosurgical units, laser surgical tools, and high-frequency ultrasonic scalpels), can severely impair endoscopic visibility, posing a notable challenge. The presence of dense smoke obstructs the surgeon’s field of view, increases the risk of surgical errors, and reduces procedural efficiency. Moreover, smoke-induced image degradation adversely affects vision-based downstream surgical tasks, such as phase recognition [Yang_2024_Surgformer_Surgical, Zou_2025_Capturing_action, Wagner_2023_Comparative_validation, Yue_2023_Cascade_Multi, Guan_2024_Label_Guided, Zhang_2026_CSAP_Assist], tool detection, semantic segmentation [Sheng_2025_AMNCutter_Affinity, Sheng_2024_Revisiting_Surgical, Yue_2024_SurgicalSAM_Efficient], depth estimation [Zou_2025_SSIFNet_Spatialtemporal, Tao_2023_SVT_SDE, Cheng_2022_Deep_Laparoscopic], 3D surface reconstruction [Gong_2024_Self_Supervised, Maier_Hein_2014_Comparative_Validation], etc., which require both high frame quality and real-time processing speed [Han_2025_Towards_Reliable, Long_2025_Surgical_embodied, Wang_2025_Learning_dissection, Xu_2026_Surgical_Action, Dou_2025_Artificial_intelligence] . At present, mechanical smoke evacuation or filtration is commonly achieved through hardware-based systems, which often require additional effort and impose extra workload during surgery, disrupting the surgical workflow and causing surgeon fatigue [Manning_2018_Laparoscopic_lens, Xia_2025_In_Vivo]. Therefore, ensuring a clear visual field in real time is critical [Ulmer_2008_The_hazards, Carbajo_Rodrguez_2009_Surgical_smoke, Weld_2007_Analysis_of].
To this end, real-time image processing has gained increasing attention as a means of digitally removing smoke during intraoperative visualization in a more streamlined and automated way by leveraging computer vision and deep learning techniques. It offers a cost-effective, low-risk, and efficient alternative for surgical smoke removal [Luo_2017_Vision_Based, Chen_2020_De_smokeGCN, Salazar_Colores_2020_Desmoking_Laparoscopy, Li_2024_Multi_frequency, Wang_2024_Desmoking_of]. However, traditional models often struggle with this task, as they were originally designed for outdoor hazy environments. Unlike natural hazy images, in vivo surgical scenes present unique challenges, including complex physics parameter estimation due to non-uniform distribution of smoke, spatially heterogeneous illumination caused by radial light attenuation, inaccurate color restoration on non-smoke objects (e.g., surgical instruments), severe information loss resulting from highly dense smoke obscuration, and temporal incoherence of smoke occurrence due to the intermittent use of energy devices [Modrzejewski_2020_Light_modelling, Azagra_2023_Endomapper_dataset, Rodrguez_Puigvert_2023_LightDepth_Single, Stelzer_2024_Generation_and, Kumar_2023_Understanding_surgical, Tian_2015_Single_Image]. To address these limitations, deep learning-based approaches have been increasingly adopted for their powerful non-linear mapping capabilities and ease of deployment in laparoscopic systems [Chen_2022_DesmokeNet_A, Li_2022_Endoscopy_image]. These methods show considerable promise for surgical desmoking, offering more consistent color restoration and improved visual clarity. Nevertheless, due to the scarcity of paired desmoking benchmarks, most deep learning-based models face the challenge of generalizability, as they can produce high-quality predictions on a dataset with synthetic smoke while suffering from performance degradation on another dataset with real smoke. This is an important concern for real-world clinical applications [Wu_2025_Self_Supervised].
To overcome these limitations, we propose the Surgical Atmospheric Model (SurgiATM), a physics-guided model tailored to the laparoscopic desmoking task. SurgiATM can be seamlessly embedded into existing neural networks with minimal modifications to their original architectures, while enhancing both performance and generalizability. Notably, SurgiATM introduces no additional trainable parameters, indicating that the observed improvements arise from the integration of physical guidance rather than merely increasing model trainable weights and complexity. Specifically, the main contributions of this study are:
  1. 1.
    We propose SurgiATM, a parameter-efficient, plug-and-play, physics-guided model, that improves performance and generalizability of surgical desmoking approaches to better meet clinical requirements.
  2. 2.
    The proposed model introduces zero trainable weights and can be seamlessly integrated into state-of-the-art (SOTA) methods built upon various backbone architectures (e.g., U-Net, Swin Transformer).
  3. 3.
    We validate SurgiATM on three public surgical datasets, demonstrating improvements in both performance and generalizability across multiple SOTA desmoking methods and various surgical scenes, thereby confirming its effectiveness, stability, and generalizability.
2 Related Work
2.1 Physics-Based Methods
The Atmospheric Scattering Model (ASM) is a widely used physical model that describes the formation of hazy natural images by modeling the absorption, scattering, and reflection of light by objects and particles in the atmospheric medium [Narasimhan_2002_Vision_and, Narasimhan_2003_Contrast_restoration]. Based on the ASM, clear images can be restored from degraded observations by estimating the required physical parameters, such as medium transmission, atmospheric light, and scene depth, which has motivated extensive research in the field. In early studies, the Gaussian distribution and Markov Random Field (MRF) model were classic and well-known techniques for estimating the parameters [Fattal_2008_Single_image, Tan_2008_Visibility_in]. Then, [Kaiming_2011_Single_Image] proposed the Dark Channel Prior (DCP) model for natural image dehazing, in which the medium transmission and global atmospheric light were statistically estimated according to their physical properties. The DCP assumes that haze in an image is uniformly distributed and that, in hazeless images, at least one color channel in most non-sky patches or pixels has very low intensity. In contrast, haze tends to raise pixel intensities across all channels, causing the dark channel to become brighter and the image to appear more whitish. Since 2012, the field of surgical desmoking has gained increasing attention, with many valuable studies published over the past decade. [Gu_2015_Virtual_Digital] developed a digital defogging system that uses the DCP to enhance image quality in laparoscopic surgery. [Kotwal_2016_Joint_desmoking] combined the desmoking and denoising tasks of laparoscopic images and transformed them into a Bayesian inference problem by proposing a novel probabilistic graphical model. [Baid_2017_Joint_desmoking] extends [Kotwal_2016_Joint_desmoking] by introducing additional image priors of color and texture derived from sparse dictionary models. In [Tchaka_2017_Chromaticity_based], histogram equalization is combined with an adaptive DCP to remove surgical smoke. [Luo_2017_Vision_Based] rearranged the original atmospheric model by introducing an atmospheric veil to represent the transmission map and then solved the parameter estimation problem via the Poisson equation in the frequency domain. [Wang_2018_Variational_based] first developed an energy function to optimize and estimate a smoke veil to recover smokeless images based on [Tian_2015_Single_Image]. Combining the hardware camera and the software algorithm, [Azam_2022_Smoke_removal] proposed a method based on multiple-exposure fusion, which enhances local contrast and color saturation, thereby improving image quality. These works offer valuable insights into the feasibility of physics-based approaches for surgical desmoking, encouraging further research.
2.2 Deep Learning-Based Methods
To achieve higher restoration quality, the powerful fitting and generative capabilities of Deep Neural Networks (DNNs) offer an alternative solution [Cai_2016_DehazeNet_An, Li_2017_AOD_Net, Dong_2020_Multi_Scale, Qin_2020_FFA_Net, Bolkar_2018_Deep_Smoke, Sidorov_2020_Generative_Smoke]. Current deep learning-based desmoking methods can be broadly categorized into two frameworks: non-adversarial learning and adversarial learning.
2.2.1 Non-adversarial Learning Methods
This framework indicates that the model is tasked with learning a mapping from smoky or hazy images to their corresponding clean versions. [Chen_2018_Unsupervised_Learning] trained a U-Net model [Ronneberger_2015_U_Net] on laparoscopic videos with synthetic smoke using Blender [^1] https://www.blender.org/. [Bolkar_2018_Deep_Smoke] attempted to address surgical smoke removal via AODNet [Li_2017_AOD_Net], a Convolutional Neural Network (CNN) designed to estimate global light illuminance and the transmission map simultaneously. Similar to [Chen_2018_Unsupervised_Learning], [Bolkar_2018_Deep_Smoke] trained the model on synthetically generated smoky in vivo frames by blending Perlin noise with clean frames. Furthermore, instead of training on synthetic smoky frames, [Ma_2021_A_Smoke] attempted to use adjacent frames, along with the presence or absence of smoke, as training inputs and corresponding ground truth. Similarly, in the most recent study [Wu_2025_Self_Supervised], they implemented a real-world self-supervised video desmoking approach by capturing the informative features from frames recorded prior to the activation of high-energy surgical instruments. Based on the above training data strategy, considerable efforts have been made to improve performance further. [Wang_2019_Multiscale_deep] designed a U-Net-like network and proposed Laplacian image pyramid decomposition to capture multi-scale features. [Chen_2020_De_smokeGCN] presented a novel generative-collaborative learning framework (De-smokeGCN), dividing the desmoking into two sub-tasks: smoke detection and smoke removal. A smoke detector predicted the smoke mask, which was leveraged as a prior for the smoke removal network and motivated several subsequent studies [Hong_2023_MARS_GAN, Zhou_2024_Synchronizing_Detection]. [Lin_2021_A_desmoking] improved the decoder part of the U-Net by appending a convolutional block attention module to generate a guidance mask for feature maps. [Kanakatte_2021_Surgical_Smoke] introduced a workflow to simultaneously remove smoke and restore colors, using U-Net as the backbone. [Ma_2024_Laparoscopic_Video] leveraged deformable convolution incorporated with a mutual attention mechanism to model temporal features and combined Local Binary Patterns with input frames as a texture prior. [Wang_2024_Smoke_veil] developed a method consisting of three U-Nets to produce distinct outputs, each serving as an element in the energy function derived from their previous study [Wang_2018_Variational_based]. [Li_2024_Multi_frequency] developed a surgical desmoking method based on the Diffusion Model [Ho_2020_Denoising_Diffusion, Jiaming_2021_Denoising_Diffusion], incorporating a multi-level frequency analysis module to integrate features across different frequency bands. In addition to convolutional network architectures, the Swin Transformer (SwinT) [Liu_2021_Swin_Transformer] has recently gained increasing attention in this field because of its hierarchical representation and efficient local-to-global attention mechanism. [Wang_2023_Surgical_smoke] proposed a network that utilized several convolutional layers to extract low-level features and then used SwinT to analyze deep and global information from smoky frames. [Wang_2024_Desmoking_of] developed a U-shaped transformer model based on U-Net and SwinT to enhance feature representation by combining local detail extraction with global context modeling.
2.2.2 Adversarial Learning Methods
Apart from conventional image-to-image translation, converting laparoscopic images from the smoky domain to the smokeless domain offers an alternative paradigm for addressing the surgical desmoking problem. By regarding desmoking as an image style transfer task, the Generative Adversarial Network (GAN) [Goodfellow_2020_Generative_adversarial] and its variants have emerged as effective solutions. [Sidorov_2020_Generative_Smoke] modeled a mapping between smoky and smoke-free domains using a conventional GAN as the backbone and leveraging the perceptual image quality score as a loss function to improve the result quality. [Salazar_Colores_2020_Desmoking_Laparoscopy] developed a method based on a Conditional Generative Adversarial Network (cGAN) [Mirza_2014_Conditional_generative] and used an embedded dark channel as a prior. Compared to the above GANs, the Cycle-Consistent Adversarial Network (CycleGAN) [Zhu_2017_Unpaired_Image] is well-suited for image style transfer tasks and is therefore more commonly used for surgical smoke removal [Su_2023_Multi_stages, Wang_2025_Desmoke_VCU, Islam_2024_LVQE_Laparoscopic]. [Vishal_2019_Guided_Unsupervised] enhanced CycleGAN by incorporating multi-scale feature extraction and introducing an upsampling loss to improve the contrast of the desmoked frames. In their subsequent studies, they further included structure-consistency loss and designed a refinement module to improve restoration quality [Vishal_2020_Unsupervised_Desmoking, Venkatesh_2020_Unsupervised_smoke]. [Pan_2022_DeSmoke_LAP] further designed internal-channel and dark-channel loss functions based on the characteristics of smoky pixels to improve model performance. Similar to [Chen_2020_De_smokeGCN], [Zhou_2024_Synchronizing_Detection] added an extra detection network to the CycleGAN architecture to estimate the smoke mask as a prior for subsequent desmoking. [Hong_2023_MARS_GAN] adopted a multilevel strategy to adaptively learn non-homogeneous smoke features, using the predicted smoke distribution as a prior.
To date, existing studies have made significant contributions; however, some limitations remain, as the feasibility and practicality of physics-based methods are primarily restricted by accuracy, while those of deep learning-based methods are mainly hindered by robustness and generalizability. In this paper, we conduct a statistical analysis of their respective advantages and subsequently derive a plug-and-play module to augment the existing end-to-end deep learning desmoking methods, further enhancing their performance and generalizability across different benchmarks.
3 Methods
The proposed method is motivated by an interesting observation in laparoscopic surgical video smoke removal: predictions from most desmoking methods yield a Laplacian-like error distribution (see Fig. 1). Building upon this finding, we introduce a new restoration formula that minimizes prediction errors and enhances model robustness by blending physics-based and deep learning-based approaches from the perspective of probability distributions.
Figure 1: The left images are from a natural dehazing dataset, O-HAZE [Ancuti_2018_O_HAZE]; the middle group is from a real-world surgical desmoking benchmark VASST-desmoke [Xia_2025_In_Vivo]; and the line chart (right) shows the difference in error distributions between surgical desmoking and natural dehazing, computed from the entire dataset. The $\text{1}^{\text{st}}$ row displays the hazy or smoky images; the $\text{2}^{\text{nd}}$ row shows the smoke or haze mask estimated from the ground truth; and the $\text{3}^{\text{rd}}$ row presents the error magnitude map of the DCP prediction.
3.1 Background
3.1.1 Physics-Based Modeling
The atmospheric scattering model [Narasimhan_2002_Vision_and, Narasimhan_2003_Contrast_restoration, Fattal_2008_Single_image, Tan_2008_Visibility_in] is typically formulated as:
$I(x,c)=J(x,c)\cdot t(x)+A(c)\cdot(1-t(x)),$ (1)
where $c$ denotes a color channel; $x$ represents the spatial location of a pixel; $I\in\mathbb{R}^{H\times W\times C}$ is the observed image, where $C$, $W$, and $H$ are the number of color channels, width and height, respectively; $J$ indicates the corresponding haze-free image; $A\in\mathbb{R}^{1\times 1\times C}$ denotes the global atmospheric light; and $t\in\mathbb{R}^{H\times W\times 1}$ is the medium transmission, which can alternatively be expressed as t=1St=1-S with smoke density $S$. In particular, $t$ and $J$ can be further formulated as:
$t(x)=e^{-\beta(\lambda)\cdot d(x)},$ (2)
$J(x,c)=A(c)\cdot\rho(x,c),$ (3)
where $\beta(\lambda)\in\mathbb{R}$ is the scattering coefficient for a specific wavelength $\lambda$ of light; $d\in\mathbb{R}^{H\times W\times 1}$ represents the scene depth; and $\rho\in\mathbb{R}^{H\times W\times C}$ is the normalized radiance [Narasimhan_2003_Contrast_restoration] or reflectance [Tan_2008_Visibility_in] of a scene point.
[Kaiming_2011_Single_Image] proposed the Dark Channel Prior (DCP), which estimates the transmission map $t$ using the dark channel:
$\begin{split}t(x)&=1-D(x)\\ &=1-\underset{c\in C}{\mathrm{min}}\left(\underset{u\in\Omega(x;z)}{\mathrm{min}}\left(\frac{I(u,c)}{A(c)}\right)\right),\end{split}$ (4)
where $D$ denotes the dark channel and $\Omega(x)$ represents a square window of size $z$ centered at location $x$. Accordingly, the restoration formula is given by:
$J_{\mathrm{DCP}}(x,c)=\frac{I(x,c)-A(c)}{1-D(x)}+A(c),$ (5)
where $J_{\mathrm{DCP}}$ indicates that the image $J$ is estimated using DCP; 1D(x)1-D(x) is empirically clipped by a lower bound $t_{0}$ in practical implementations to prevent division by zero, while this application-based adjustment is omitted for simplicity and clarity of the subsequent mathematical derivation.
3.1.2 Deep Learning-Based Modeling
Clearly, in pixels where the smoke density is high and white or gray surgical instruments are present (i.e., $D\rightarrow 1$), the conventional physics-based model (5) may fail. Moreover, radial light attenuation can violate the assumption that $A\in\mathbb{R}^{1\times 1\times C}$; under such conditions, $A$ is more appropriately represented as $A\in\mathbb{R}^{H\times W\times C}$. Adhering to the original assumption may introduce more errors, whereas adopting the latter formulation significantly increases the degrees of freedom of the solution space. Therefore, Deep Neural Network (DNN), a powerful technique for non-linear mapping, is widely employed for smoke removal. Its formulation and training objective are typically expressed as:
[…]
$\theta^{*}=\underset{\theta}{\arg\min}\;\;$ $\mathcal{L}(J,J_{\mathrm{DNN}};\mathcal{X},\theta),$ (7)
where $\Phi$ indicates a non-linear mapping parameterized by trainable weights $\theta$; $\theta^{*}$ represents the optimized weights obtained through training on the dataset $\mathcal{X}$; $\tilde{Z}$ and $\mathcal{L}$ denote priors and loss function, respectively, as determined by the specific method design. Nevertheless, its performance and generalizability may be limited when deployed in unseen surgical scenarios.
3.2 Mixture-of-Experts and Optimization
Inspired by the Mixture-of-Experts (MoE), an effective mechanism for complementary advantages [Jacobs_1991_Adaptive_Mixtures, Ding_2025_DenseFormer_MoE], we begin with an ideal gating model:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=\tilde{W}(x)\cdot J_{\mathrm{DCP}}(x,c)\\ &\quad+(1-\tilde{W}(x))\cdot J_{\mathrm{DNN}}(x,c),\end{split}$ (8)
where $J_{\mathrm{MoE}}$ is the restored result of the gating model, and $\tilde{W}\in\{0,1\}^{H\times W\times 1}$ is a pixel-wise indicator specifying which method is activated for smoke removal. Specifically, $\tilde{W}(x)=1$ if $J_{\mathrm{DCP}}(x,c)$ is deemed more reliable than $J_{\mathrm{DNN}}(x,c)$; otherwise, $\tilde{W}(x)=0$. Optimizing $\tilde{W}$ facilitates the most accurate prediction. For simplification and tractability, $\tilde{W}$ is relaxed to a continuous domain, yielding $W\in[0,1]^{H\times W\times 1}$, and the restorations are formulated using additive error models:
$J_{\mathrm{DCP}}(x,c)$ $=J(x,c)+\varepsilon_{\mathrm{DCP}}(x,c),$ (9)
$J_{\mathrm{DNN}}(x,c)$ $=J(x,c)+\varepsilon_{\mathrm{DNN}}(x,c),$ (10)
where the predicted image is represented as the ground truth augmented by an independent and identically distributed (i.i.d.) error term $\varepsilon\in\mathbb{R}^{H\times W\times C}$. The error term is commonly assumed to follow a zero-mean Gaussian distribution with variance $\sigma^{2}$, denoted as $\varepsilon(x,c)\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(\mu=0,\sigma^{2})$, which is a classic assumption in statistical learning. Then, we have:
$J_{\mathrm{MoE}}(x,c)$ $=J(x,c)+\varepsilon_{\mathrm{MoE}}(x,c),$ (11)
$\varepsilon_{\mathrm{MoE}}(x,c)$ $=W(x)\cdot\varepsilon_{\mathrm{DCP}}(x,c)$
$\quad+(1-W(x))\cdot\varepsilon_{\mathrm{DNN}}(x,c).$ (12)
Clearly, optimizing $J_{\mathrm{MoE}}$ is equivalent to minimizing the error term $\varepsilon_{\mathrm{MoE}}$:
$\begin{split}\min_{W(x)}\;\;&\mathbb{E}[\varepsilon_{\mathrm{MoE}}^{2}],\quad s.t.\;\;0\leq W(x)\leq 1.\end{split}$ (13)
Table 1: Comparison of Prediction Error Distributions with Gaussian and Laplacian Distributions via Jensen-Shannon (JS) Divergence in VASST-desmoke with real-world smoke and Cholec80 with synthetic smoke
Method VASST-desmoke Cholec80
Gaussian Laplacian Gaussian Laplacian
DCP [Kaiming_2011_Single_Image] 0.044989 0.013664 0.002782 0.010886
AODNet [Li_2017_AOD_Net] 0.077348 0.056151 0.006405 0.012263
CGAN-DC [Salazar_Colores_2020_Desmoking_Laparoscopy] 0.033480 0.033894 0.007939 0.005278
De-smokeGCN [Chen_2020_De_smokeGCN] 0.027138 0.001988 0.007484 0.005802
GCANet [Chen_2019_Gated_Context] 0.009448 0.003621 0.003164 0.008474
LGUTransformer [Wang_2024_Desmoking_of] 0.010483 0.003129 0.016931 0.003623
MARS-GAN [Hong_2023_MARS_GAN] 0.024441 0.064570 0.009878 0.005369
MSBDN [Dong_2020_Multi_Scale] 0.020062 0.007690 0.005995 0.005536
RSTN [Wang_2023_Surgical_smoke] 0.026959 0.006900 0.012523 0.004193
SSIM-PAN [Sidorov_2020_Generative_Smoke] 0.018315 0.011999 0.005900 0.006517
SVPNet [Wang_2024_Smoke_veil] 0.022240 0.000675 0.007876 0.006492
Avg. 0.028628 0.013795 0.008734 0.006734
Figure 2: The Laplacian parameters (i.e., $\mu$ and $b$) estimated from the specific dataset, and indicator $W^{*}$ calculated with (19), are presented across different methods with the Confidence Interval (CI). The upper and lower rows correspond to Cholec80 and VASTT-desmoke, respectively.
Interestingly, in the context of laparoscopic surgical video smoke removal, we observe that, for real-world surgical desmoking, the error terms do not follow a Gaussian-like distribution; rather, they follow a Laplacian-like distribution, as shown in Fig. 1 and reported in Table 1, where most methods yield lower Jenson-Shannon (JS) divergence to Laplacian distribution. We attribute this phenomenon to the non-uniform and temporally incoherent nature of real-world surgical smoke, which, compared to natural haze, results in more zero-error regions in smokeless areas and frames, with notable errors at smoky pixels. Based on this finding, we first model the error terms in surgical desmoking as zero-mean Laplacian, that is, $\varepsilon\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Laplace}\left(\mu=0,b\right)$, where $\mu$ and $b$ are the location and scale parameters, respectively. Furthermore, our analysis indicates that $\mu$ and $b$ are potentially correlated with the dark channel $D$ (see Fig. 2). Hence, in the proposed approach, dark channel $D$ is incorporated as a conditioning factor for the non-zero-mean Laplacian distribution:
$\varepsilon_{\mathcal{M}}(x,c)\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Laplace}\left(\mu_{\mathcal{M}}(D(x)),b_{\mathcal{M}}(D(x))\right),$ (14)
where $\mathcal{M}$ represents a method, either DCP or DNN. Collecting (12), and (14), we have:
$\begin{split}\mathrm{Var}[\varepsilon_{\mathrm{MoE}}]&=2\cdot\big(W^{2}(x)\cdot b^{2}_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))^{2}\cdot b^{2}_{\mathrm{DNN}}(D(x))\big),\end{split}$ (15)
$\begin{split}\mathbb{E}[\varepsilon_{\mathrm{MoE}}]&=W(x)\cdot\mu_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))\cdot\mu_{\mathrm{DNN}}(D(x)).\end{split}$ (16)
Then, the original optimization problem (13) is rewritten as:
$\min_{W(x)}\;\;$ $\mathrm{Var}(\varepsilon_{\mathrm{MoE}})+\left(\mathbb{E}(\varepsilon_{\mathrm{MoE}})\right)^{2},$ (17)
$s.t.\;\;$ $0\leq W(x)\leq 1.$ (18)
We solve (17) and obtain the unconstrained solution:
$W^{*}(x;D)=\frac{2b^{2}_{\mathrm{2},x}-\mu_{\mathrm{2},x}\cdot(\mu_{\mathrm{1},x}-\mu_{\mathrm{2},x})}{2b^{2}_{\mathrm{1},x}+2b^{2}_{\mathrm{2},x}+(\mu_{\mathrm{1},x}-\mu_{\mathrm{2},x})^{2}},$ (19)
where $W^{*}(x;D)$ is the optimized $W(x)$ conditioned on the dark channel $D$ without constraints. For clarity, $b_{\mathrm{DCP}}(D(x))$ and $\mu_{\mathrm{DNN}}(D(x))$ are written as $b_{1,x}$ and $\mu_{2,x}$, respectively. Subsequently, $W^{*}(x)$ is clipped to the interval $[0,1]$ to satisfy the constraint (18).
3.3 A New Restoration Formula
Following (19), Fig. 2 presents the trends of $W^{*}$ given $D$, based on statistics across all methods in two datasets: Cholec80 [Twinanda_2017_EndoNet_A] with synthetic smoke, and the benchmark VASST-desmoke [Xia_2024_A_New, Xia_2025_In_Vivo] containing real-world smoke. Deep learning models display varied trends, while a consistent negative correlation between $W^{*}$ and $D$ is observed. Therefore, in this work, $W$ is estimated by:
$\hat{W}(x)=1-D(x),$ (20)
where $\hat{W}$ represents the approximated $W$ given $D$. The motivations are as follows:
  1. 1.
    Developing a general plug-and-play mechanism for the existing DNN methods is one of the objectives of this study, and this function captures the overall correlation between $W$ and $D$;
  2. 2.
    Equation (20) is intuitive: a high intensity of dark channel (i.e., $D(x)\rightarrow 1$) statistically indicates dense smoke [Kaiming_2011_Single_Image, Xia_2025_In_Vivo], which challenges DCP due to information loss and division by zero or very small values, thereby making DNN the dominant model (i.e., $W(x)\rightarrow 0$);
  3. 3.
    It encourages consistent predictions in smokeless regions (i.e., $D(x)\rightarrow 0$);
  4. 4.
    The denominator 1D1-D and the unknown parameter $A$ in (5) are eliminated in the subsequent derivation, stabilizing the computation, simplifying the modeling, and precluding hard value clipping.
By combining (5), (8), and (20), we obtain:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot J_{\mathrm{DNN}}(x,c),\end{split}$ (21)
where $A$ remains an unknown parameter that requires an additional model for its estimation. To address this issue, we first use (3) to further decompose $J_{\mathrm{DNN}}$:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot A(c)\cdot\rho_{\mathrm{DNN}}(x,c),\end{split}$ (22)
where $\rho_{\mathrm{DNN}}(x,c)=J_{\mathrm{DNN}}(x,c)/A(c)$ represents the reflectance calculated from the normalized DNN prediction. Equation (22) can be reformulated as:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=I(x,c)-D(x)\cdot A(c)\cdot(1-\rho_{\mathrm{DNN}}(x,c))\\ &=I(x,c)-\mathcal{D}(x)\cdot(1-\rho_{\mathrm{DNN}}(x,c)),\end{split}$ (23)
where $\mathcal{D}$ denotes the denormalized dark channel $D$. Recalling (4), $\mathcal{D}$ is defined as:
$\mathcal{D}(x)=D(x)\cdot A(c)=\underset{c\in C}{\mathrm{min}}\left(\underset{u\in\Omega(x;z)}{\mathrm{min}}\left(I(u,c)\right)\right),$ (24)
where $A$ is eliminated, thereby avoiding the challenging estimation of global light illuminance in laparoscopic surgical scenes. In this case, the original DNN modeling in (6) and (7) can be improved as:
[…]
$\theta^{*}=\underset{\theta}{\arg\min}\;\;$ $\mathcal{L}\left(J,(I-\mathcal{D}\cdot(1-\rho_{\mathrm{DNN}}));\mathcal{X},\theta\right),$ (26)
where $\sigma$ is a differentiable function (typically a sigmoid) that maps the DNN output to the interval $[0,1]$. It can be ignored if the output is already constrained within this range, depending on the specific model design. The proposed formulation introduces only minimal modifications to the original DNN model $\Phi_{\theta}$, without adding any additional trainable parameters. Notably, SurgiATM leverages the denormalized dark channel as a key element, in line with recent methods [Chen_2019_Gated_Context, Hong_2023_MARS_GAN, Chen_2020_De_smokeGCN], suggesting that the dark channel prior plays a useful role in surgical desmoking [Xia_2025_In_Vivo].
3.4 Gradient Computation and Refinement
In this section, we compute the gradient with respect to the model prediction $\varrho_{\mathrm{DNN}}$ for two commonly used loss functions: Mean Absolute Error (MAE or L1) loss and Mean Squared Error (MSE or L2) loss. Their gradients are denoted as $\nabla_{\varrho}\mathcal{L}_{1}$ and $\nabla_{\varrho}\mathcal{L}_{2}$, respectively:
$\nabla_{\varrho}\mathcal{L}_{1}=\begin{cases}-\mathcal{D},&J>1-\mathcal{D}\cdot(1-\rho);\\ \mathcal{D},&J<1-\mathcal{D}\cdot(1-\rho);\\ 0,&J=1-\mathcal{D}\cdot(1-\rho),\end{cases}$ (27)
$\nabla_{\varrho}\mathcal{L}_{2}=-\mathcal{D}\cdot(J-(1-\mathcal{D}\cdot(1-\rho))),$ (28)
where, for simplicity, $\rho_{\mathrm{DNN}}$ is written as $\rho$. A potential issue arises when $\mathcal{D}(x)=0$, as this results in a zero gradient, preventing the DNN model from updating its weights at those pixels. To address this, we refine $\mathcal{D}$ as:
$\hat{\mathcal{D}}=\frac{\mathcal{D}+\eta}{1+\eta},\quad\eta\in[0,+\infty),$ (29)
where $\eta$ is a smoothing factor; $\eta=0$ corresponds to no refinement; and $\eta\rightarrow+\infty$ implies $\hat{\mathcal{D}}\rightarrow 1$, causing (26) to degenerate into conventional residual prediction. Ablation results for $\eta$ and $z$ are reported in Section 4.5.
4 Experiments And Results
4.1 Implementation Details
By default, we set the smoothing factor $\eta=0.1$ and the window size $z=15$ [Kaiming_2011_Single_Image]. All experiments are conducted on a 24GB NVIDIA RTX 3090 GPU. To ensure fair comparisons, training configurations are largely standardized across all methods, employing the Adam optimizer with a learning rate of 0.0002 and epochs of 50. Due to GPU memory constraints, the batch size depends on their respective publications.
4.2 Datasets and Metric
4.2.1 Datasets
In this study, we utilize three datasets: Cholec80 [Twinanda_2017_EndoNet_A], VASST-desmoke [Xia_2024_A_New, Xia_2025_In_Vivo], and the Hamlyn Centre Laparoscopic and Endoscopic Dataset Videos (referred to as the Hamlyn Dataset for simplicity) [Giannarou_2013_Probabilistic_Tracking, Ye_2017_Self_Supervised]. The Hamlyn Dataset is specifically reserved for external evaluation. For Cholec80, 500 smokeless and 500 smoky frames are randomly sampled from each video [Leibetseder_2017_Real_Time], with the smoke-free frames then blended with synthetic smoke [Bolkar_2018_Deep_Smoke] for training. VASST-desmoke is a recently released paired benchmark for laparoscopic surgical smoke removal. It provides both real smoky frames and corresponding ground-truth smoke-free annotations, thereby enabling evaluation on real surgical smoke with intuitive full-reference metrics. Potentially, VASST-desmoke also allows models to learn from real smoke. The Hamlyn Dataset is excluded from all training and statistical analysis procedures in this work, and its frames are sampled from the videos at 25 frames per second (FPS). We perform five-fold cross-evaluation, ensuring no video overlap between the training and evaluation sets. The three datasets record various surgical procedures with inconsistent camera devices and lighting environments in different hospitals, significantly challenging the stability and generalizability of surgical desmoking methods.
4.2.2 Metrics
For the evaluation of paired datasets, we leverage full-reference metrics: CIEDE2000, PSNR, RMSE, and SSIM, to evaluate perceptual color difference, reconstruction quality, prediction error, and structure similarity, respectively. For the unpaired datasets, we evaluate model performance using non-reference metrics, including BRISQUE and NIQE to measure how close an image is to natural scenes, and FADE [Choi_2015_Referenceless_Prediction], which specifically measures fog-related degeneration.
Table 2: Training and Evaluation on VASST-desmoke (Baseline / Baseline + SurgiATM)
Method Year Backbone Loss Function CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet [Li_2017_AOD_Net] 2017 CNN MSE 10.349 / 6.327 17.378 / 21.994 0.147 / 0.090 0.693 / 0.789
GCANet [Chen_2019_Gated_Context] 2019 CNN MSE 7.893 / 5.369 21.091 / 23.907 0.095 / 0.073 0.729 / 0.806
De-smokeGCN [Chen_2020_De_smokeGCN] 2020 U-Net MAE + Others 9.577 / 7.348 19.779 / 22.127 0.115 / 0.094 0.550 / 0.717
MSBDN [Dong_2020_Multi_Scale] 2020 U-Net MSE 6.679 / 5.164 21.604 / 23.881 0.088 / 0.073 0.651 / 0.813
SVPNet [Wang_2024_Smoke_veil] 2024 U-Net MSE + Others 5.574 / 5.141 23.464 / 24.362 0.073 / 0.070 0.789 / 0.820
CGAN-DC [Chen_2019_Gated_Context] 2019 cGAN Adversarial + Others 6.735 / 5.812 22.453 / 23.809 0.083 / 0.079 0.676 / 0.782
MARS-GAN [Hong_2023_MARS_GAN] 2023 CycleGAN Adversarial + Others 6.389 / 5.899 20.121 / 23.488 0.113 / 0.076 0.602 / 0.819
SSIM-PAN [Sidorov_2020_Generative_Smoke] 2020 GAN Adversarial + Others 7.046 / 5.811 21.465 / 23.477 0.090 / 0.075 0.655 / 0.750
RSTN [Wang_2023_Surgical_smoke] 2023 SwinT Charbonnier + Others 5.385 / 4.976 23.263 / 24.306 0.074 / 0.071 0.750 / 0.824
LGUTransformer [Wang_2024_Desmoking_of] 2024 SwinT + U-Net Charbonnier 9.927 / 5.990 18.238 / 23.376 0.121 / 0.080 0.510 / 0.761
Table 3: Training on Cholec80 with synthetic smoke and evaluation on VASST-desmoke
Method CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet [Li_2017_AOD_Net] 9.609 / 7.774 18.621 / 21.526 0.132 / 0.109 0.729 / 0.798
GCANet [Chen_2019_Gated_Context] 9.803 / 7.042 19.338 / 22.220 0.119 / 0.097 0.747 / 0.801
De-smokeGCN [Chen_2020_De_smokeGCN] 8.370 / 7.585 20.654 / 22.088 0.109 / 0.102 0.716 / 0.788
MSBDN [Dong_2020_Multi_Scale] 8.277 / 7.327 20.360 / 22.153 0.114 / 0.103 0.769 / 0.810
SVPNet [Wang_2024_Smoke_veil] 8.105 / 7.797 21.196 / 21.712 0.110 / 0.109 0.806 / 0.793
CGAN-DC [Chen_2019_Gated_Context] 8.177 / 7.525 20.663 / 21.988 0.113 / 0.105 0.770 / 0.805
MARS-GAN [Hong_2023_MARS_GAN] 7.808 / 7.444 21.444 / 22.110 0.107 / 0.104 0.792 / 0.811
SSIM-PAN [Sidorov_2020_Generative_Smoke] 8.350 / 7.600 20.338 / 21.805 0.115 / 0.106 0.761 / 0.806
RSTN [Wang_2023_Surgical_smoke] 7.963 / 7.533 21.034 / 22.065 0.111 / 0.107 0.763 / 0.813
LGUTransformer [Wang_2024_Desmoking_of] 7.357 / 7.352 22.108 / 22.111 0.105 / 0.104 0.812 / 0.818
Table 4: Training with synthetic smoke and evaluation with real smoke on Cholec80
Method BRISQUE $\downarrow$ FADE $\downarrow$ NIQE $\downarrow$
AODNet [Li_2017_AOD_Net] 12.551 / 14.814 0.440 / 0.484 10.449 / 9.434
GCANet [Chen_2019_Gated_Context] 7.795 / 11.566 0.407 / 0.403 9.866 / 9.331
De-smokeGCN [Chen_2020_De_smokeGCN] 4.857 / 5.575 0.384 / 0.376 10.885 / 9.569
MSBDN [Dong_2020_Multi_Scale] 11.838 / 12.176 0.421 / 0.411 10.299 / 9.026
SVPNet [Wang_2024_Smoke_veil] 12.812 / 12.134 0.416 / 0.402 10.579 / 9.387
CGAN-DC [Chen_2019_Gated_Context] 10.258 / 9.945 0.413 / 0.404 9.517 / 9.039
MARS-GAN [Hong_2023_MARS_GAN] 13.999 / 11.680 0.416 / 0.413 13.336 / 13.884
SSIM-PAN [Sidorov_2020_Generative_Smoke] 10.539 / 11.093 0.411 / 0.403 9.811 / 11.999
RSTN [Wang_2023_Surgical_smoke] 13.868 / 12.652 0.420 / 0.414 12.245 / 16.910
LGUTransformer [Wang_2024_Desmoking_of] 9.126 / 11.236 0.417 / 0.416 13.734 / 12.471
Table 5: Training on VASST-desmoke and evaluation on Hamlyn Dataset
Method BRISQUE $\downarrow$ FADE $\downarrow$ NIQE $\downarrow$
AODNet [Li_2017_AOD_Net] 36.549 / 30.515 0.331 / 0.322 12.418 / 11.851
GCANet [Chen_2019_Gated_Context] 13.354 / 26.555 0.324 / 0.294 11.512 / 10.983
De-smokeGCN [Chen_2020_De_smokeGCN] 33.236 / 28.007 0.329 / 0.325 29.448 / 18.802
MSBDN [Dong_2020_Multi_Scale] 26.992 / 25.198 0.384 / 0.322 14.224 / 10.944
SVPNet [Wang_2024_Smoke_veil] 39.862 / 30.325 0.401 / 0.317 16.245 / 13.306
CGAN-DC [Chen_2019_Gated_Context] 14.052 / 16.699 0.352 / 0.336 16.616 / 13.216
MARS-GAN [Hong_2023_MARS_GAN] 33.722 / 30.601 0.377 / 0.369 15.011 / 16.381
SSIM-PAN [Sidorov_2020_Generative_Smoke] 11.936 / 16.047 0.357 / 0.298 16.613 / 20.297
RSTN [Wang_2023_Surgical_smoke] 35.629 / 30.232 0.412 / 0.330 16.442 / 12.645
LGUTransformer [Wang_2024_Desmoking_of] 35.260 / 29.706 0.295 / 0.294 15.300 / 12.135
Table 6: Training on Cholec80 with synthetic smoke and evaluation on Hamlyn Dataset
Method BRISQUE $\downarrow$ FADE $\downarrow$ NIQE $\downarrow$
AODNet [Li_2017_AOD_Net] 30.292 / 31.168 0.426 / 0.432 11.074 / 11.834
GCANet [Chen_2019_Gated_Context] 19.366 / 26.102 0.385 / 0.293 11.628 / 11.083
De-smokeGCN [Chen_2020_De_smokeGCN] 13.920 / 17.647 0.330 / 0.315 11.628 / 11.467
MSBDN [Dong_2020_Multi_Scale] 26.630 / 27.559 0.409 / 0.332 12.152 / 11.882
SVPNet [Wang_2024_Smoke_veil] 29.065 / 28.375 0.406 / 0.267 14.358 / 12.292
CGAN-DC [Chen_2019_Gated_Context] 25.175 / 24.127 0.408 / 0.315 10.445 / 10.741
MARS-GAN [Hong_2023_MARS_GAN] 29.811 / 27.287 0.409 / 0.311 16.443 / 15.136
SSIM-PAN [Sidorov_2020_Generative_Smoke] 24.605 / 26.001 0.395 / 0.312 10.852 / 19.337
RSTN [Wang_2023_Surgical_smoke] 26.305 / 28.347 0.414 / 0.328 13.875 / 14.217
LGUTransformer [Wang_2024_Desmoking_of] 23.974 / 26.497 0.383 / 0.346 16.786 / 17.448
4.3 Comparison of Quantitative Results
We train five folds of ten existing methods and their corresponding SurgiATM modified versions, using Cholec80 with synthetic smoke and VASST-desmoke with real-world smoke, respectively. We then compare their five-fold average performance on the real-world surgical smoke datasets: VASST-desmoke, Cholec80 (real smoke), and the Hamlyn Dataset. In total, we train 200 models and conduct 500 evaluations.
The quantitative results on VASST-desmoke are presented in Tables 2 and 3. They demonstrate that incorporating SurgiATM effectively improves model performance, regardless of the backbone architecture or loss function. However, for SVPNet in Table 3, SurgiATM yields no improvement in SSIM. We argue that, compared to other desmoking methods, SVPNet [Wang_2024_Smoke_veil] employs an energy function [Wang_2018_Variational_based] for restoration. This is a physics-guided approach similar to ours, and its three physics-related parameters are predicted by three independent U-Nets. As a result, incorporating an extra physics-guided component (our SurgiATM) into SVPNet may slightly disrupt the internal physical consistency, leading to the observed deterioration in SSIM.
For Cholec80, we report the evaluation results in Table 4. Our SurgiATM effectively improves most model performances in terms of the fog-related metric (i.e., FADE score). In terms of the external evaluation on the Hamlyn Dataset, as shown in Tables 6 and 5, for the defogging metric FADE, employing our SurgiATM enhances most methods, except for the traditional dehazing method AODNet in Table 6. We attribute this to two factors: 1) the simple architecture and insufficient trainable weights of AODNet (i.e., only five convolutional layers) compared to other methods, and 2) the non-Laplacian error distribution of AODNet when trained on Cholec80 with synthetic smoke, as reported in Table 1. In contrast, although GCANet and SSIM-PAN also exhibit non-Laplacian error distributions in Cholec80, and CGAN-DC and MARS-GAN in VASST-desmoke, their performances are effectively enhanced owing to both their well-designed architectures and the strong generalizability of SurgiATM.
On the other hand, SurgiATM demonstrates limited effectiveness in terms of BRISQUE and NIQE across the ten existing methods in Tables 4, 5, and 6. The two possible reasons are: 1) these two metrics are designed to measure similarity to natural images, which may not generalize well to surgical endoscopic frames due to the substantial domain gap between them [Wang_2024_Desmoking_of]; and 2) the Laplacian-guided error minimization of SurgiATM may contribute to dissimilarity from natural images (Gaussian-like error).
Figure 3: Error comparison in VASST-desmoke. “+” indicates the baseline integrated with our SurgiATM. ”GT” represents ground truth. The methods are trained in Cholec80 with synthetic smoke. Orange boxes mark the errors declined by SurgiATM. Overall, “Baseline + SurgiATM” exhibits more accurate restoration.
Figure 4: Comparison of SurgiATM with baselines on real datasets: Cholec80 (left two columns) and the Hamlyn Dataset (right two columns). Blue boxes highlight the comparatively accurate and stable restoration achieved by SurgiATM.
4.4 Comparison of Qualitative Results
In this section, we present qualitative comparisons across several representative baselines: De-smokeGCN, GCANet, MARS-GAN, and SVPNet, each selected for different reasons. Specifically, De-smokeGCN [Chen_2020_De_smokeGCN] is a classic DNN-based desmoking method introduced in 2020; GCANet [Chen_2019_Gated_Context] is designed for natural dehazing and deraining; MARS-GAN represents an adversarial-learning-based approach; and SVPNet [Wang_2024_Smoke_veil], published in 2024, uses a physics-based loss function tailored for surgical desmoking. We train the models on Cholec80 with synthetic smoke and evaluate them on VASST-desmoke, Cholec80, and the Hamlyn Dataset with real surgical smoke, challenging their stability and generalizability.
As shown in the error maps in Fig. 3, SurgiATM effectively reduces desmoking errors in a computationally efficient and interpretable manner, making it well-suited for deployment on edge devices and for stable surgical desmoking. For the natural dehazing method GCANet, it causes color distoration in Fig. 3(a), which is corrected to a certain degree by using SurgiATM. In addition, the proposed approach can mitigate the grid artifact (see the first two rows in Fig. 3), demonstrating promising stability. As shown in Fig. 4, the desmoking results of the baselines are improved to varying degrees. For instance, De-smokeGCN and GCANet can produce more stable results with SurgiATM by reducing the grid artifacts (see Fig. 4(a)-(d) of De-smokeGCN) and correcting the restoration colors (see Fig. 4(d) of GCANet). Moreover, SurgiATM helps baselines eliminate more smoke, as shown in Fig. 4(a) for MARS-GAN and Fig. 4(c) for SVPNet, GCANet, and MARS-GAN.
4.5 Ablation Study
We conduct ablation studies to investigate the influence of two hyperparameters (i.e., smoothing factor $\eta$ and shifting window size $z$) on model performance. Experiments are carried out on the real-world benchmark VASST-desmoke using three representative baselines (i.e., AODNet, RSTN, and SSIM-PAN), chosen for their minimal reliance on auxiliary modules and loss functions, thereby enabling a clear assessment of the impact of the only two hyperparameters, smoother $\eta$ and window size $z$, with minimal confounding factors. The RMSE metrics are averaged from five-fold cross-validation and reported in Fig. 5. We observed that for most methods, $\eta=0.1$ usually leads to better performance. No obvious trend is observed for the window size $z$: a larger window size is preferred for AODNet and SSIM-PAN, while a smaller one is better for RSTN. Therefore, for the use of SurgiATM, $\eta=0.1$ is a recommended configuration, whereas the optimal window size should be determined for each specific method.
Figure 5: Each heatmap illustrates the ablation study for a specific baseline model incorporated with SurgiATM, showing the impact of different hyperparameter settings in terms of the RMSE metric (lower is better). Each heatmap uses an independent color scale.
5 Conclusion and Future Work
In this work, we propose SurgiATM for laparoscopic surgical smoke removal, which can be seamlessly incorporated into existing DNN-based desmoking methods without additional trainable parameters. By using SurgiATM, the performance and generalizability of these methods are enhanced, enabling the generation of cleaner frames for both surgeons and downstream tasks. We analyze the discrepancy between natural dehazing and surgical desmoking and design SurgiATM by first formulating a Mixture-of-Experts model. We then conduct a comprehensive statistical analysis of existing desmoking methods and derive a new restoration formula tailored for surgical smoke removal. Extensive experiments across three distinct datasets validate the effectiveness of SurgiATM.
Nevertheless, several limitations accompany its advantages. First, the indicator $W$ is currently estimated as a linear function of the dark channel $D$, which captures the general trends across methods, and thus leaves some room for improvement when focusing on a particular method; this could potentially be addressed by deriving method-specific formulations. Second, the denormalized dark channel obtained from DCP remains relatively coarse-grained. Developing an appropriate refinement strategy may further improve its performance and result quality; for example, using guided image filtering or attention mechanisms could generate a comparatively fine-grained denormalized dark channel.