### v2

# SurgiATM: A Physics-Guided Plug-and-Play Model for Deep Learning-Based Smoke Removal in Laparoscopic Surgery

Abstract
During laparoscopic surgery, smoke generated by tissue cauterization can significantly degrade the visual quality of endoscopic frames, increasing the risk of surgical errors and hindering both clinical decision-making and computer-assisted visual analysis. Consequently, removing surgical smoke is critical to ensuring patient safety and maintaining operative efficiency. In this study, we propose the Surgical Atmospheric Model (SurgiATM) for surgical smoke removal. SurgiATM statistically bridges a physics-based atmospheric model and data-driven deep learning models, combining the superior generalizability of the former with the high accuracy of the latter. Furthermore, SurgiATM is designed as a lightweight, plug-and-play module that can be seamlessly integrated into diverse surgical desmoking architectures to enhance their accuracy and stability. The proposed method is derived via statistically optimizing Mixture-of-Experts (MoE) model at the output end of arbitrary deep learning methods, with a Laplacian-like error distribution specifically leveraged to model surgical smoke. The output-stage MoE ensures minimal modification to the architecture of the original methods, while the Laplacian-like distribution characteristic of surgical smoke enables a lightweight reconstruction formulation with minimal parameters. Therefore, SurgiATM introduces only two hyperparameters and no additional trainable weights, preserving the original network architecture with minimal computational and modification overhead. We conduct extensive experiments on three public surgical datasets with ten desmoking methods, involving multiple network architectures and covering diverse procedures, including cholecystectomy, partial nephrectomy, and diaphragm dissection. The results demonstrate that incorporating SurgiATM commonly reduces the restoration errors of existing models and relatively enhances their generalizability, without adding any trainable layers or weights. This highlights the convenience, low cost, effectiveness, and generalizability of the proposed method. The code for SurgiATM is released at https://github.com/MingyuShengSMY/SurgiATM.
keywords:
Endoscopy , Laparoscopic surgery , Smoke removal , Plug-and-play , Parameter-efficient
journal: Medical Image Analysis
[…]
[usyd]organization=School of Computer Science, The University of Sydney,city=Sydney, state=NSW, country=Australia
[…]
[sjt]organization=Institute of Medical Robotics, School of Biomedical Engineering, Shanghai Jiao Tong University,city=Shanghai, country=China
[…]
[hvd]organization=Department of Radiology, Brigham and Women’s Hospital, and Harvard Medical School,city=Boston, state=MA, country=United States
1 Introduction
Minimally invasive surgery (MIS) has noticeably benefited patients by reducing risk and enabling faster recovery (Maier-Hein et al., 2017, 2022). However, surgical smoke, commonly generated by high-energy instruments (e.g., electrosurgical units, laser surgical tools, and high-frequency ultrasonic scalpels), can severely impair endoscopic visibility, posing a notable challenge (Duelmer et al., 2026). The presence of dense smoke obstructs the surgeon’s field of view, increases the risk of surgical errors, and reduces procedural efficiency. Moreover, smoke-induced image degradation adversely affects vision-based downstream surgical tasks, such as phase recognition (Yang et al., 2024; Zou et al., 2025a; Wagner et al., 2023; Yue et al., 2023; Guan et al., 2024; Zhang et al., 2025; Das et al., 2025), tool detection, semantic segmentation (Sheng et al., 2025, 2024; Yue et al., 2024; Sestini et al., 2025; Cartucho et al., 2024), depth estimation (Zou et al., 2025b; Tao et al., 2023; Cheng et al., 2022), 3D surface reconstruction (Gong et al., 2024; Maier-Hein et al., 2014; Rau et al., 2024), etc., which require both high frame quality and real-time processing speed (Han et al., 2025; Long et al., 2025; Wang et al., 2025a; Xu et al., 2026; Dou et al., 2025; Rueckert et al., 2026). At present, mechanical smoke evacuation or filtration is commonly achieved through hardware-based systems, which often require additional effort and impose extra workload during surgery, disrupting the surgical workflow and causing surgeon fatigue (Manning et al., 2018; Xia et al., 2025). Therefore, ensuring a clear visual field in real time is critical (Ulmer, 2008; Carbajo-Rodríguez et al., 2009; Weld et al., 2007).
To this end, real-time image processing has gained increasing attention as a means of digitally removing smoke during intraoperative visualization in a more streamlined and automated way by leveraging computer vision and deep learning techniques. It offers a cost-effective, low-risk, and efficient alternative for surgical smoke removal (Luo et al., 2017; Chen et al., 2020; Salazar-Colores et al., 2020; Li et al., 2024; Wang et al., 2024b). However, traditional models often struggle with this task, as they were originally designed for outdoor hazy environments. Unlike natural hazy images, in vivo surgical scenes present unique challenges, including complex physics parameter estimation due to non-uniform distribution of smoke, spatially heterogeneous illumination caused by radial light attenuation, inaccurate color restoration on non-smoke objects (e.g., surgical instruments), severe information loss resulting from highly dense smoke obscuration, and temporal incoherence of smoke occurrence due to the intermittent use of energy devices (Modrzejewski et al., 2020; Azagra et al., 2023; Rodríguez-Puigvert et al., 2023; Stelzer et al., 2024; Kumar et al., 2023; Tian et al., 2015; Ramesh et al., 2023). To address these limitations, deep learning-based approaches have been increasingly adopted for their powerful non-linear mapping capabilities and ease of deployment in laparoscopic systems (Chen et al., 2022; Li et al., 2022). These methods show considerable promise for surgical desmoking, offering more consistent color restoration and improved visual clarity. Nevertheless, due to the scarcity of paired desmoking benchmarks, most deep learning-based models face the challenge of generalizability, as they can produce high-quality predictions on a dataset with synthetic smoke while suffering from performance degradation on another dataset with real smoke. This is an important concern for real-world clinical applications (Wu et al., 2025b).
To overcome these limitations, we propose the Surgical Atmospheric Model (SurgiATM), a physics-guided model tailored to the laparoscopic desmoking task. SurgiATM can be seamlessly embedded into existing neural networks with minimal modifications to their original architectures, while enhancing both performance and generalizability. Notably, SurgiATM introduces no additional trainable parameters, indicating that the observed improvements arise from the integration of physical guidance rather than merely increasing model trainable weights and complexity. Specifically, the main contributions of this study are:
  • We propose SurgiATM, a parameter-efficient, plug-and-play, physics-guided model, that improves performance and generalizability of surgical desmoking approaches, providing clearer desmoking results.
  • The proposed model is derived via Mixture-of-Experts (MoE) model at the output end, and thus SurgiATM can be employed as an output layer (or module), seamlessly integrated into state-of-the-art (SOTA) methods built upon various backbone architectures (e.g., U-Net, Swin Transformer).
  • We develop SurgiATM based on statistical analysis and physical model; thereby it introduces zero trainable weights, while only two hyper-parameters are involved. The recommended configurations for the two hyper-parameters are analyzed in ablation study.
  • We validate SurgiATM on three public surgical datasets, demonstrating improvements in both performance and generalizability across multiple SOTA desmoking methods and various laparoscopic surgical scenes, thereby confirming its effectiveness, low efficiency overhead, stability, and generalizability for laparoscopic surgical smoke removal.
2 Related Work
2.1 Physics-Based Methods
The Atmospheric Scattering Model (ASM) is a widely used physical model that describes the formation of hazy natural images by modeling the absorption, scattering, and reflection of light by objects and particles in the atmospheric medium (Narasimhan and Nayar, 2002, 2003). Based on the ASM, clear images can be restored from degraded observations by estimating the required physical parameters, such as medium transmission, atmospheric light, and scene depth, which has motivated extensive research in the field. In early studies, the Gaussian distribution and Markov Random Field (MRF) model were classic and well-known techniques for estimating the parameters (Fattal, 2008; Tan, 2008). Then, Kaiming et al. (2011) proposed the Dark Channel Prior (DCP) model for natural image dehazing, in which the medium transmission and global atmospheric light were statistically estimated according to their physical properties. The DCP assumes that haze in an image is uniformly distributed and that, in hazeless images, at least one color channel in most non-sky patches or pixels has very low intensity. In contrast, haze tends to raise pixel intensities across all channels, causing the dark channel to become brighter and the image to appear more whitish. Since 2012, the field of surgical desmoking has gained increasing attention, with many valuable studies published over the past decade. Gu et al. (2015) developed a digital defogging system that uses the DCP to enhance image quality in laparoscopic surgery. Kotwal et al. (2016) combined the desmoking and denoising tasks of laparoscopic images and transformed them into a Bayesian inference problem by proposing a novel probabilistic graphical model. Baid et al. (2017) extends (Kotwal et al., 2016) by introducing additional image priors of color and texture derived from sparse dictionary models. In Tchaka et al. (2017) study, histogram equalization is combined with an adaptive DCP to remove surgical smoke. Luo et al. (2017) rearranged the original atmospheric model by introducing an atmospheric veil to represent the transmission map and then solved the parameter estimation problem via the Poisson equation in the frequency domain. Wang et al. (2018) first developed an energy function to optimize and estimate a smoke veil to recover smokeless images based on Tian et al. (2015). Combining the hardware camera and the software algorithm, Azam et al. (2022) proposed a method based on multiple-exposure fusion, which enhances local contrast and color saturation, thereby improving image quality. These works offer valuable insights into the feasibility of physics-based approaches for surgical desmoking, encouraging further research.
2.2 Deep Learning-Based Methods
To achieve higher restoration quality, the powerful fitting and generative capabilities of Deep Neural Networks (DNNs) offer an alternative solution (Cai et al., 2016; Li et al., 2017; Dong et al., 2020; Qin et al., 2020; Bolkar et al., 2018; Sidorov et al., 2020). Current deep learning-based desmoking methods can be broadly categorized into two frameworks: non-adversarial learning and adversarial learning.
2.2.1 Non-adversarial Learning Methods
This framework indicates that the model is tasked with learning a mapping from smoky or hazy images to their corresponding clean versions. Chen et al. (2018) trained a U-Net model (Ronneberger et al., 2015) on laparoscopic videos with synthetic smoke using Blender [^1] https://www.blender.org/. Bolkar et al. (2018) attempted to address surgical smoke removal via AODNet (Li et al., 2017), a Convolutional Neural Network (CNN) designed to estimate global light illuminance and the transmission map simultaneously. Similar to Chen et al. (2018), Bolkar et al. (2018) trained the model on synthetically generated smoky in vivo frames by blending Perlin noise with clean frames. Furthermore, instead of training on synthetic smoky frames, Ma et al. (2021) attempted to use adjacent frames, along with the presence or absence of smoke, as training inputs and corresponding ground truth. Similarly, in the most recent study, Wu et al. (2025b) implemented a real-world self-supervised video desmoking approach by capturing the informative features from frames recorded prior to the activation of high-energy surgical instruments. Based on the above training data strategy, considerable efforts have been made to improve performance further. Wang et al. (2019) designed a U-Net-like network and proposed Laplacian image pyramid decomposition to capture multi-scale features. Chen et al. (2020) presented a novel generative-collaborative learning framework (De-smokeGCN), dividing the desmoking into two sub-tasks: smoke detection and smoke removal. A smoke detector predicted the smoke mask, which was leveraged as a prior for the smoke removal network and motivated several subsequent studies (Hong et al., 2023; Zhou et al., 2024). Lin et al. (2021) improved the decoder part of the U-Net by appending a convolutional block attention module to generate a guidance mask for feature maps. Kanakatte et al. (2021) introduced a workflow to simultaneously remove smoke and restore colors, using U-Net as the backbone. Ma et al. (2024a) leveraged deformable convolution incorporated with a mutual attention mechanism to model temporal features and combined Local Binary Patterns with input frames as a texture prior. Wang et al. (2024a) developed a method consisting of three U-Nets to produce distinct outputs, each serving as an element in the energy function derived from their previous study (Wang et al., 2018). Li et al. (2024) developed a surgical desmoking method based on the Diffusion Model (Ho et al., 2020; Jiaming et al., 2021), incorporating a multi-level frequency analysis module to integrate features across different frequency bands. Wu et al. (2025a) further combine the information in frequency and temporal domains, and developed a multi-scale Gaussian window constraint to capture inter-frame features with fine structural details preserved. In addition to convolutional network architectures, the Swin Transformer (SwinT) (Liu et al., 2021) has recently gained increasing attention in this field because of its hierarchical representation and efficient local-to-global attention mechanism. Wang et al. (2023) proposed a network that utilized several convolutional layers to extract low-level features and then used SwinT to analyze deep and global information from smoky frames. Wang et al. (2024b) developed a U-shaped transformer model based on U-Net and SwinT to enhance feature representation by combining local detail extraction with global context modeling.
Figure 1: The proposed SurgiATM is derived from Atmospheric Scattering Model, Normalized Radiance Decomposition, Mixture-of-Experts (MoE), and Laplacian Error Minimization. We collect prediction results and errors from a lot of SOTA methods with diverse backbone architectures; and then Laplacian error distribution is observed from surgical desmoking which is different from natural fog. Compared to physics-based methods (e.g., DCP), data-driven methods usually demonstrate higher accuracy, while the formers show strong stability. To complement their strengths, SurgiATM is developed by optimizing this system.
2.2.2 Adversarial Learning Methods
Apart from conventional image-to-image translation, converting laparoscopic images from the smoky domain to the smokeless domain offers an alternative paradigm for addressing the surgical desmoking problem. By regarding desmoking as an image style transfer task, the Generative Adversarial Network (GAN) (Goodfellow et al., 2020) and its variants have emerged as effective solutions. Sidorov et al. (2020) modeled a mapping between smoky and smoke-free domains using a conventional GAN as the backbone and leveraging the perceptual image quality score as a loss function to improve the result quality. Salazar-Colores et al. (2020) developed a method based on a Conditional Generative Adversarial Network (cGAN) (Mirza and Osindero, 2014) and used an embedded dark channel as a prior. Compared to the above GANs, the Cycle-Consistent Adversarial Network (CycleGAN) (Zhu et al., 2017) is well-suited for image style transfer tasks and is therefore more commonly used for surgical smoke removal (Su and Wu, 2023; Wang et al., 2025b; Islam et al., 2024). Vishal et al. (2019) enhanced CycleGAN by incorporating multi-scale feature extraction and introducing an upsampling loss to improve the contrast of the desmoked frames. In their subsequent studies, they further included structure-consistency loss and designed a refinement module to improve restoration quality (Vishal et al., 2020; Venkatesh et al., 2020). Pan et al. (2022) further designed internal-channel and dark-channel loss functions based on the characteristics of smoky pixels to improve model performance. Similar to Chen et al. (2020), Zhou et al. (2024) added an extra detection network to the CycleGAN architecture to estimate the smoke mask as a prior for subsequent desmoking. Hong et al. (2023) adopted a multilevel strategy to adaptively learn non-homogeneous smoke features, using the predicted smoke distribution as a prior.
To date, existing studies have made significant contributions; however, some limitations remain, as the feasibility and practicality of physics-based methods are primarily restricted by accuracy, while those of deep learning-based methods are mainly hindered by robustness and generalizability. In this paper, we conduct a statistical analysis of their respective advantages and subsequently derive a plug-and-play module to augment the existing end-to-end deep learning desmoking methods, further enhancing their performance and generalizability across different benchmarks.
Table 1: Primary Notations.
Variable Definition
II Smoky image
$J$ Smoke-free image
$_{\text{DCP}},J_{\text{DNN}} JJ$ predicted by DCP or DNN
$t$ Medium transmission
AA Global atmospheric light
$\rho$ Normalized radiance
$D$ Dark channel
$\mathcal{D}$ Denormalized dark channel
$W$ Expert-weights for MoE
$\varepsilon$ Error term
$\mu,b$ Laplacian parameters
$\eta$ Smoothing factor
$z$ Window size
3 Methods
The proposed method is motivated by an intriguing observation in laparoscopic surgical video smoke removal: predictions from most desmoking methods yield a Laplacian-like error distribution (see Fig. 2). Building upon this insight, we develop a new restoration formula that minimizes prediction errors and enhances model robustness by revisiting classic physical model, analyzing tge predictions of deep learning-based approaches, leveraging a MoE model, and optimizing from the perspective of probability distributions (see Fig. 1). SurgiATM can be seamlessly embedded at the output stage of deep learning methods, where it predicts a normalized radiance used for reconstruction. Integration with general DNN baseline methods is demonstrated in Fig. 4. The primary variables are summarized in Table 1.
Figure 2: The left images are from a natural dehazing dataset, O-HAZE (Ancuti et al., 2018); the middle group is from a real-world surgical desmoking benchmark VASST-desmoke (Xia et al., 2025); and the line chart (right) shows the difference in error distributions between surgical desmoking and natural dehazing, computed from the entire dataset. The $\text{1}^{\text{st}}$ row displays the hazy or smoky images; the $\text{2}^{\text{nd}}$ row shows the smoke or haze mask estimated from the ground truth; and the $\text{3}^{\text{rd}}$ row presents the error magnitude map of the DCP prediction.
3.1 Background
3.1.1 Physics-Based Modeling
The atmospheric scattering model (Narasimhan and Nayar, 2002, 2003; Fattal, 2008; Tan, 2008) is typically formulated as:
$\ I(x,c)=J(x,c)\cdot t(x)+A(c)\cdot(1-t(x)),$ (1)
where $c$ denotes a color channel; $x$ represents the spatial location of a pixel; $I\in\mathbb{R}^{H\times W\times C}$ is the observed image, where $C$, $W$, and $H$ are the number of color channels, width and height, respectively; $J$ indicates the corresponding haze-free image; $A\in\mathbb{R}^{1\times 1\times C}$ denotes the global atmospheric light; and $t\in\mathbb{R}^{H\times W\times 1}$ is the medium transmission. In particular, $t$ and $J$ can be further formulated as:
$\ t(x)=e^{-\beta(\lambda)\cdot d(x)},$ (2)
$\ J(x,c)=A(c)\cdot\rho(x,c),$ (3)
where $\beta(\lambda)\in\mathbb{R}$ is the scattering coefficient for a specific wavelength $\lambda$ of light; $d\in\mathbb{R}^{H\times W\times 1}$ represents the scene depth; and $\rho\in\mathbb{R}^{H\times W\times C}$ is the normalized radiance of a scene point, which is mainly determined by reflectance of object surfaces (Narasimhan and Nayar, 2003; Tan, 2008).
Kaiming et al. (2011) proposed the Dark Channel Prior (DCP), which estimates the transmission map $t$ using the dark channel:
$\ \begin{aligned} t(x)&=1-D(x)\\ &=1-\underset{c\in C}{\mathrm{min}}\left(\underset{u\in\Omega(x;z)}{\mathrm{min}}\left(\frac{I(u,c)}{A(c)}\right)\right),\end{aligned}$ (4)
where $D$ denotes the dark channel and $\Omega(x)$ represents a square window of size $z$ centered at location $x$. Accordingly, the restoration formula is given by:
$\ J_{\mathrm{DCP}}(x,c)=\frac{I(x,c)-A(c)}{1-D(x)}+A(c),$ (5)
where $J_{\mathrm{DCP}}$ indicates that the image $J$ is estimated using DCP; 1D(x)1-D(x) is empirically clipped by a lower bound $t_{0}$ in practical implementations to prevent division by zero, while this application-based adjustment is omitted for simplicity and clarity of the subsequent mathematical derivation.
3.1.2 Deep Learning-Based Modeling
Clearly, in pixels where the smoke density is high and white or gray surgical instruments are present (i.e., $D\rightarrow 1$), the conventional physics-based model Eq. (5) may fail. Moreover, radial light attenuation can violate the assumption that $A\in\mathbb{R}^{1\times 1\times C}$; under such conditions, $A$ is more appropriately represented as $A\in\mathbb{R}^{H\times W\times C}$. Adhering to the original assumption may introduce more errors, whereas adopting the latter formulation significantly increases the degrees of freedom of the solution space. Therefore, Deep Neural Network (DNN), a powerful technique for non-linear mapping, is widely employed for smoke removal. Its formulation and training objective are typically expressed as:
[…]
$\ \theta^{*}=\underset{\theta}{\arg\min}\;\;\mathcal{L}(J,J_{\mathrm{DNN}};\mathcal{X},\theta),$ (7)
where $\Phi$ indicates a non-linear mapping parameterized by trainable weights $\theta$; $\theta^{*}$ represents the optimized weights obtained through training on the dataset $\mathcal{X}$; $\tilde{Z}$ and $\mathcal{L}$ denote priors and loss function, respectively, as determined by the specific method design. Nevertheless, its generalizability may be limited when deployed in unseen surgical scenarios.
3.2 Mixture-of-Experts and Optimization
Inspired by the Mixture-of-Experts (MoE), an effective mechanism for complementary advantages (Jacobs et al., 1991; Ding et al., 2025), we begin with an ideal gating model:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=\tilde{W}(x)\cdot J_{\mathrm{DCP}}(x,c)\\ &\quad+(1-\tilde{W}(x))\cdot J_{\mathrm{DNN}}(x,c),\end{aligned}$ (8)
where $J_{\mathrm{MoE}}$ is the restored result of the gating model, and $\tilde{W}\in\{0,1\}^{H\times W\times 1}$ is a pixel-wise indicator specifying which method is activated for smoke removal. Specifically, $\tilde{W}(x)=1$ if $J_{\mathrm{DCP}}(x,c)$ is deemed more reliable than $J_{\mathrm{DNN}}(x,c)$; otherwise, $\tilde{W}(x)=0$. Optimizing $\tilde{W}$ facilitates the most accurate prediction. For simplification and tractability, $\tilde{W}$ is relaxed to a continuous domain, yielding $W\in[0,1]^{H\times W\times 1}$, and the restorations are formulated using additive error models:
$\ J_{\mathrm{DCP}}(x,c)=J(x,c)+\varepsilon_{\mathrm{DCP}}(x,c),$ (9)
$\ J_{\mathrm{DNN}}(x,c)=J(x,c)+\varepsilon_{\mathrm{DNN}}(x,c),$ (10)
where the predicted image is represented as the ground truth augmented by an independent and identically distributed (i.i.d.) error term $\varepsilon\in\mathbb{R}^{H\times W\times C}$. The error term is commonly assumed to follow a zero-mean Gaussian distribution with variance $\sigma^{2}$, denoted as $\varepsilon(x,c)\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(\mu=0,\sigma^{2})$, which is a classic assumption in statistical learning. Then, we have:
$\ J_{\mathrm{MoE}}(x,c)=J(x,c)+\varepsilon_{\mathrm{MoE}}(x,c),$ (11)
$\ \begin{aligned} \varepsilon_{\mathrm{MoE}}(x,c)=W(x)\cdot\varepsilon_{\mathrm{DCP}}(x,c)\\ \quad+(1-W(x))\cdot\varepsilon_{\mathrm{DNN}}(x,c).\end{aligned}$ (12)
Clearly, optimizing $J_{\mathrm{MoE}}$ is equivalent to minimizing the error term $\varepsilon_{\mathrm{MoE}}$:
$\ \begin{aligned} \min_{W(x)}\;\;&\mathbb{E}[\varepsilon_{\mathrm{MoE}}^{2}],\quad s.t.\;\;0\leq W(x)\leq 1.\end{aligned}$ (13)
Table 2: Comparison of Prediction Error Distributions with Gaussian and Laplacian Distributions via Jensen-Shannon (JS) Divergence in VASST-desmoke with real-world smoke and Cholec80 with synthetic smoke.
Method VASST-desmoke Cholec80
Gaussian Laplacian Gaussian Laplacian
DCP (Kaiming et al., 2011) 0.044989 0.013664 0.002782 0.010886
AODNet (Li et al., 2017) 0.077348 0.056151 0.006405 0.012263
CGAN-DC (Salazar-Colores et al., 2020) 0.033480 0.033894 0.007939 0.005278
De-smokeGCN (Chen et al., 2020) 0.027138 0.001988 0.007484 0.005802
GCANet (Chen et al., 2019) 0.009448 0.003621 0.003164 0.008474
LGUTransformer (Wang et al., 2024b) 0.010483 0.003129 0.016931 0.003623
MARS-GAN (Hong et al., 2023) 0.024441 0.064570 0.009878 0.005369
MSBDN (Dong et al., 2020) 0.020062 0.007690 0.005995 0.005536
RSTN (Wang et al., 2023) 0.026959 0.006900 0.012523 0.004193
SSIM-PAN (Sidorov et al., 2020) 0.018315 0.011999 0.005900 0.006517
SVPNet (Wang et al., 2024a) 0.022240 0.000675 0.007876 0.006492
Avg. 0.028628 0.013795 0.008734 0.006734
Figure 3: The Laplacian parameters (i.e., $\mu$ and $b$) estimated from the specific dataset, and indicator $W^{*}$ calculated with Eq. (19), are presented across different methods with the Confidence Interval (CI). The upper and lower rows correspond to Cholec80 and VASTT-desmoke, respectively.
Interestingly, in the context of laparoscopic surgical video smoke removal, we observe that, for real-world surgical desmoking, the error terms do not follow a Gaussian-like distribution; rather, they follow a Laplacian-like distribution, as shown in Fig. 2 and reported in Table 2, where most methods yield lower Jenson-Shannon (JS) divergence to Laplacian distribution. We attribute this phenomenon to the non-uniform and temporally incoherent nature of real-world surgical smoke, which, compared to natural haze, results in more zero-error regions in smokeless areas and frames, with notable errors at smoky pixels. Based on this finding, we first model the error terms in surgical desmoking as zero-mean Laplacian, that is, $\varepsilon\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Laplace}\left(\mu=0,b\right)$, where $\mu$ and $b$ are the location and scale parameters, respectively. Furthermore, our analysis indicates that $\mu$ and $b$ are potentially correlated with the dark channel $D$ (see Fig. 3). Hence, in the proposed approach, dark channel $D$ is incorporated as a conditioning factor for the non-zero-mean Laplacian distribution:
$\ \varepsilon_{\mathcal{M}}(x,c)\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Laplace}\left(\mu_{\mathcal{M}}(D(x)),b_{\mathcal{M}}(D(x))\right),$ (14)
where $\mathcal{M}$ represents a method, either DCP or DNN. Collecting Eqs.(12), and (14), we have:
$\ \begin{aligned} \mathrm{Var}[\varepsilon_{\mathrm{MoE}}]&=2\cdot\big(W^{2}(x)\cdot b^{2}_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))^{2}\cdot b^{2}_{\mathrm{DNN}}(D(x))\big),\end{aligned}$ (15)
$\ \begin{aligned} \mathbb{E}[\varepsilon_{\mathrm{MoE}}]&=W(x)\cdot\mu_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))\cdot\mu_{\mathrm{DNN}}(D(x)).\end{aligned}$ (16)
Then, the original optimization problem Eq. (13) is rewritten as:
$\ \min_{W(x)}\;\;\mathrm{Var}(\varepsilon_{\mathrm{MoE}})+\left(\mathbb{E}(\varepsilon_{\mathrm{MoE}})\right)^{2},$ (17)
$\ s.t.\;\;0\leq W(x)\leq 1.$ (18)
We solve Eq. (17) and obtain the unconstrained solution:
$\ W^{*}(x;D)=\frac{2b^{2}_{\mathrm{2},x}-\mu_{\mathrm{2},x}\cdot(\mu_{\mathrm{1},x}-\mu_{\mathrm{2},x})}{2b^{2}_{\mathrm{1},x}+2b^{2}_{\mathrm{2},x}+(\mu_{\mathrm{1},x}-\mu_{\mathrm{2},x})^{2}},$ (19)
where $W^{*}(x;D)$ is the optimized $W(x)$ conditioned on the dark channel $D$ without constraints. For simplicity, $b_{\mathrm{DCP}}(D(x))$, $b_{\mathrm{DNN}}(D(x))$, $\mu_{\mathrm{DCP}}(D(x))$ and $\mu_{\mathrm{DNN}}(D(x))$ are written as $b_{1,x}$, $b_{2,x}$, $\mu_{1,x}$ and $\mu_{2,x}$, respectively. Subsequently, $W^{*}(x)$ is clipped to the interval $[0,1]$ to satisfy the constraint Eq. (18).
3.3 A New Restoration Formula
Following Eq. (19), Fig. 3 presents the trends of $W^{*}$ given $D$, based on statistics across all methods in two datasets: Cholec80 (Twinanda et al., 2017) with synthetic smoke, and the benchmark VASST-desmoke (Xia et al., 2024, 2025) containing real-world smoke. Deep learning models display varied trends, while a consistent negative correlation between $W^{*}$ and $D$ is observed. Therefore, in this work, $W$ is estimated by:
$\ \hat{W}(x)=1-D(x),$ (20)
Table 3: Pearson Correlation Coefficient between $W$ and $D$.
Method VASST-desmoke Cholec80
AODNet (Li et al., 2017) -0.065 -0.962
CGAN-DC (Salazar-Colores et al., 2020) -0.820 -0.963
De-smokeGCN (Chen et al., 2020) -0.763 -0.973
GCANet (Chen et al., 2019) -0.933 -0.765
LGUTransformer (Wang et al., 2024b) -0.899 -0.954
MARS-GAN (Hong et al., 2023) -0.977 -0.942
MSBDN (Dong et al., 2020) -0.899 -0.964
RSTN (Wang et al., 2023) -0.896 -0.945
SSIM-PAN (Sidorov et al., 2020) -0.842 -0.982
SVPNet (Wang et al., 2024a) -0.873 -0.965
Avg. -0.797 -0.942
where $\hat{W}$ represents the approximated $W$ given $D$. The motivations are as follows:
  1. 1.
    We calculate Pearson Correlation Coefficient (PCC) between $W$ and $D$ for all methods. Most PCCs are close to 0.9-0.9, that strong negative linear correlation is obsereved, as reported in Table 3. For the low PCC of AODNet in VASST-desmoke, we observe that the origin AODNet struggles to deal with dense smoke (i.e., $D$ $\rightarrow$ 1), showing similar quantitative performance to DCP (i.e., $W$ $\rightarrow$ 0.5 means equal significance for DCP and AODNet); this is because AODNet is featured with only five convolutional layers and exhibits weaker reconstruction ability facing to dense smoke and fewer training data compared to Cholec80.
  2. 2.
    Developing a general plug-and-play mechanism for the existing DNN methods is one of the objectives of this study, and this function captures the overall correlation between $W$ and $D$;
  3. 3.
    Equation Eq. (20) is intuitive: a high intensity of dark channel (i.e., $D(x)\rightarrow 1$) statistically indicates dense smoke (Kaiming et al., 2011; Xia et al., 2025), which challenges DCP due to information loss and division by zero or very small values, thereby making DNN the dominant model (i.e., $W(x)\rightarrow 0$);
  4. 4.
    It encourages consistent predictions in smokeless regions (i.e., $D(x)\rightarrow 0$);
  5. 5.
    The denominator 1D1-D and the unknown parameter $A$ in Eq. (5) are eliminated in the subsequent derivation, stabilizing the computation, simplifying the modeling, and precluding hard value clipping.
By combining Eqs.(5), (8), and (20), we obtain:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot J_{\mathrm{DNN}}(x,c).\end{aligned}$ (21)
Then, we use the physical reflection model Eq. (3) to further decompose $J_{\mathrm{DNN}}$:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot A(c)\cdot\rho_{\mathrm{DNN}}(x,c),\end{aligned}$ (22)
where $\rho_{\mathrm{DNN}}(x,c)=J_{\mathrm{DNN}}(x,c)/A(c)$ represents the normalized radiance obtained from Eq. (3). Equation Eq. (22) can be reformulated as:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=I(x,c)-D(x)\cdot A(c)\cdot(1-\rho_{\mathrm{DNN}}(x,c))\\ &=I(x,c)-\mathcal{D}(x)\cdot(1-\rho_{\mathrm{DNN}}(x,c)),\end{aligned}$ (23)
where $\mathcal{D}$ denotes the denormalized dark channel $D$. Recalling Eq. (4), $\mathcal{D}$ is defined as:
$\ \mathcal{D}(x)=D(x)\cdot A(c)=\underset{c\in C}{\mathrm{min}}\left(\underset{u\in\Omega(x;z)}{\mathrm{min}}\left(I(u,c)\right)\right),$ (24)
where $A$ is eliminated, thereby avoiding the challenging estimation of global light illuminance in laparoscopic surgical scenes. In this case, the original DNN modeling in Eqs.(6) and (7) can be improved as:
[…]
$\ \theta^{*}=\underset{\theta}{\arg\min}\;\;\mathcal{L}\left(J,(I-\mathcal{D}\cdot(1-\rho_{\mathrm{DNN}}));\mathcal{X},\theta\right),$ (26)
where $\sigma$ is a differentiable function (typically a sigmoid) that maps the DNN output to the interval $[0,1]$. It can be ignored if the output is already constrained within this range, depending on the specific model design. The proposed formulation introduces only minimal modifications to the original DNN model $\Phi_{\theta}$, without adding any additional trainable parameters. Notably, SurgiATM leverages the denormalized dark channel as a key element, in line with recent methods (Chen et al., 2019; Hong et al., 2023; Chen et al., 2020), suggesting that the dark channel prior plays a useful role in surgical desmoking (Xia et al., 2025).
3.4 Gradient Computation and Refinement
In this section, we compute the gradient with respect to the model prediction $\rho_{\mathrm{DNN}}$ for two commonly used loss functions: Mean Absolute Error (MAE or L1) loss and Mean Squared Error (MSE or L2) loss. Their gradients are denoted as $\nabla_{\rho}\mathcal{L}_{1}$ and $\nabla_{\rho}\mathcal{L}_{2}$, respectively:
$\ \nabla_{\rho}\mathcal{L}_{1}=\begin{cases}-\mathcal{D},&J>1-\mathcal{D}\cdot(1-\rho);\\ \mathcal{D},&J<1-\mathcal{D}\cdot(1-\rho);\\ 0,&J=1-\mathcal{D}\cdot(1-\rho),\end{cases}$ (27)
$\ \nabla_{\rho}\mathcal{L}_{2}=-\mathcal{D}\cdot(J-(1-\mathcal{D}\cdot(1-\rho))),$ (28)
where, for simplicity, $\rho_{\mathrm{DNN}}$ is written as $\rho$. A potential issue arises when $\mathcal{D}(x)=0$, as this results in a zero gradient, preventing the DNN model from updating its weights at those pixels. To address this, we refine $\mathcal{D}$ as:
$\ \hat{\mathcal{D}}=\frac{\mathcal{D}+\eta}{1+\eta},\quad\eta\in[0,+\infty),$ (29)
where $\eta$ is a smoothing factor; $\eta=0$ corresponds to no refinement; and $\eta\rightarrow+\infty$ implies $\hat{\mathcal{D}}\rightarrow 1$, causing Eq. (26) to degenerate into conventional residual prediction. Ablation results for $\eta$ and $z$ are reported in Section 5.1.
Figure 4: (a): General baseline framework for DNN-based desmoking methods, and the dashed box represents the part of discriminator for GAN-based methods. (b): General framework for using SurgiATM. The modification does not involve main architecture (i.e., DNN models), suggesting minimal adjustment overhead.
3.5 Plug into Baselines
SurgiATM can be regarded as a processing pipeline (or output layer) appended to the end of DNN models (see Fig. 4, Fig. 4). It introduces only minimal modification to the original baseline. In conventional settings, the output of a DNN model is directly treated as desmoked results, based on which losses are computed and gradients are propagated during training (see Fig. 4(a)).
With SurgiATM, the DNN output is first interpreted as a normalized radiance $\rho_{\text{DNN}}$, and then passed into SurgiATM. Losses are computed using the reconstructed results produced by SurgiATM, and gradients are backpropagated through SurgiATM to the baseline DNN models. For GAN-based approaches, after incorporating SurgiATM, the discriminator should no longer operate directly on the final output of generator; instead, it should take the SurgiATM-processed results as input (see Fig. 4(b)). In particular, for GAN-based baselines that pass intermediate feature maps from the generator to the discriminator, SurgiATM should not be applied to these intermediate representations or used as a shortcut, as it is specifically designed to accept normalized radiance with physical meaning.
On the other hand, Diffusion-based methods have recently emerged as a prevalent paradigm in generative AI, and have also been explored for the surgical desmoking task. However, SurgiATM does not currently support diffusion-based approaches. This is because SurgiATM is developed for methods that directly predict smokeless images in a single forward pass and is derived from classic physical models that formulate a direct mapping from smoky frames to smokeless ones. This formulation is fundamentally different from Diffusion models, in which DNNs are charged to predicting Gaussian noise components to be removed (i.e., learning a mapping from noisy images to Gaussian noise) through an iterative denoising process involving multiple inference steps. In addition, comparatively higher computational cost of Diffusion models remains a challenge for real-time deployment (Kazerouni et al., 2023). Consequently, Diffusion-based methods are beyond the scope of this work.
Table 4: Training and evaluation on VASST-desmoke (real smoke). There is no overlap videos between training and evaluation sets. The results are represented in the form of ”Baseline / Baseline + SurgiATM”.
Method Backbone Loss Function CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet (Li et al., 2017) CNN MSE 10.349 / 6.327 17.378 / 21.994 0.147 / 0.090 0.693 / 0.789
GCANet (Chen et al., 2019) CNN MSE 7.893 / 5.369 21.091 / 23.907 0.095 / 0.073 0.729 / 0.806
De-smokeGCN (Chen et al., 2020) U-Net MAE + Others 9.577 / 7.348 19.779 / 22.127 0.115 / 0.094 0.550 / 0.717
MSBDN (Dong et al., 2020) U-Net MSE 6.679 / 5.164 21.604 / 23.881 0.088 / 0.073 0.651 / 0.813
SVPNet (Wang et al., 2024a) U-Net MSE + Others 5.574 / 5.141 23.464 / 24.362 0.073 / 0.070 0.789 / 0.820
CGAN-DC (Chen et al., 2019) cGAN Adversarial + Others 6.735 / 5.812 22.453 / 23.809 0.083 / 0.079 0.676 / 0.782
MARS-GAN (Hong et al., 2023) CycleGAN Adversarial + Others 6.389 / 5.899 20.121 / 23.488 0.113 / 0.076 0.602 / 0.819
SSIM-PAN (Sidorov et al., 2020) GAN Adversarial + Others 7.046 / 5.811 21.465 / 23.477 0.090 / 0.075 0.655 / 0.750
RSTN (Wang et al., 2023) SwinT Charbonnier + Others 5.385 / 4.976 23.263 / 24.306 0.074 / 0.071 0.750 / 0.824
LGUTransformer (Wang et al., 2024b) SwinT + U-Net Charbonnier 9.927 / 5.990 18.238 / 23.376 0.121 / 0.080 0.510 / 0.761
4 Experiments and Results
4.1 Implementation Details
For all comparisons, we set the smoothing factor $\eta=0.1$ because this value contributes to better results in out ablation study; and set the window size $z=15$ inherited from DCP (Kaiming et al., 2011). The frames are resized into $\times256$. All experiments are conducted on a 24GB NVIDIA RTX 3090 GPU. SurgiATM is extended as a output layer at the end of all approaches, without any modification to original architectures, input forms, loss functions and hyper-parameters. To ensure fair comparisons, training configurations are standardized across all methods, employing the Adam optimizer with a learning rate of 0.0002, beta1 of 0.9, beta2 of 0.999, and epochs of 50. Due to GPU memory constraints, the batch size depends on their respective publications. Please visit to https://github.com/MingyuShengSMY/SurgiATM for more implementation details. We implement SurgiATM as a torch.nn.Module and upload it onto the link, which can be directly used.
4.2 Datasets and Metric
4.2.1 Datasets
In this study, we utilize three datasets: Cholec80 (Twinanda et al., 2017), VASST-desmoke (Xia et al., 2024, 2025), and the Hamlyn Centre Laparoscopic and Endoscopic Dataset Videos (referred to as the Hamlyn Dataset for simplicity) (Giannarou et al., 2013; Ye et al., 2017). The Hamlyn Dataset is specifically reserved for external evaluation. For Cholec80, 500 smokeless and 500 smoky frames are randomly sampled from each video (Leibetseder et al., 2017), with the smoke-free frames then blended with synthetic smoke (Bolkar et al., 2018) for training. VASST-desmoke is a recently released paired benchmark for laparoscopic surgical smoke removal. It provides both real smoky frames and corresponding ground-truth smoke-free annotations, thereby enabling evaluation on real surgical smoke with intuitive full-reference metrics. Potentially, VASST-desmoke also allows models to learn from real smoke. In addition, Xia et al. (2025) further expand the VASST-desmoke, producing more paired frames sourced from Cholec80. For convenience, we note the latest part of the dataset as ”VASST-desmoke-C”. The Hamlyn Dataset is excluded from all training and statistical analysis procedures in this work, and its frames are sampled from the videos at 25 frames per second (FPS). We perform five-fold cross-evaluation, ensuring no video overlap between the training and evaluation sets. The three datasets record various surgical procedures with inconsistent camera devices and lighting environments in different hospitals, significantly challenging the stability and generalizability of surgical desmoking methods.
4.2.2 Metrics
For the evaluation of paired datasets, we leverage full-reference metrics: CIEDE2000, PSNR, RMSE, and SSIM, to evaluate perceptual color difference, reconstruction quality, prediction error, and structure similarity, respectively. For the unpaired datasets, we evaluate model performance using non-reference metrics, including BRISQUE and NIQE to measure how close an image is to natural scenes, and FADE (Choi et al., 2015), which specifically measures fog-related degeneration.
Table 5: Training and evaluation on VASST-desmoke-C (real smoke).
Method CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet (Li et al., 2017) 7.344 / 5.297 20.727 / 24.655 0.098 / 0.065 0.860 / 0.900
GCANet (Chen et al., 2019) 7.453 / 4.291 22.566 / 26.426 0.080 / 0.053 0.812 / 0.901
De-smokeGCN (Chen et al., 2020) 6.558 / 5.118 24.340 / 26.125 0.067 / 0.059 0.725 / 0.837
MSBDN (Dong et al., 2020) 4.861 / 3.999 25.532 / 27.496 0.058 / 0.051 0.834 / 0.905
SVPNet (Wang et al., 2024a) 4.509 / 3.743 26.663 / 28.121 0.052 / 0.047 0.892 / 0.915
CGAN-DC (Chen et al., 2019) 5.746 / 4.599 25.429 / 26.969 0.059 / 0.054 0.782 / 0.832
MARS-GAN (Hong et al., 2023) 4.863 / 4.592 26.096 / 27.276 0.055 / 0.051 0.888 / 0.910
SSIM-PAN (Sidorov et al., 2020) 5.382 / 4.282 25.092 / 27.003 0.061 / 0.052 0.799 / 0.877
RSTN (Wang et al., 2023) 4.321 / 3.532 26.361 / 28.395 0.053 / 0.047 0.854 / 0.915
LGUTransformer (Wang et al., 2024b) 16.862 / 5.538 17.866 / 27.532 0.130 / 0.075 0.189 / 0.895
Table 6: Training on Cholec80 (synthetic smoke) and evaluation on VASST-desmoke (real smoke).
Method CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet (Li et al., 2017) 9.609 / 7.774 18.621 / 21.526 0.132 / 0.109 0.729 / 0.798
GCANet (Chen et al., 2019) 9.803 / 7.042 19.338 / 22.220 0.119 / 0.097 0.747 / 0.801
De-smokeGCN (Chen et al., 2020) 8.370 / 7.585 20.654 / 22.088 0.109 / 0.102 0.716 / 0.788
MSBDN (Dong et al., 2020) 8.277 / 7.327 20.360 / 22.153 0.114 / 0.103 0.769 / 0.810
SVPNet (Wang et al., 2024a) 8.105 / 7.797 21.196 / 21.712 0.110 / 0.109 0.806 / 0.793
CGAN-DC (Chen et al., 2019) 8.177 / 7.525 20.663 / 21.988 0.113 / 0.105 0.770 / 0.805
MARS-GAN (Hong et al., 2023) 7.808 / 7.444 21.444 / 22.110 0.107 / 0.104 0.792 / 0.811
SSIM-PAN (Sidorov et al., 2020) 8.350 / 7.600 20.338 / 21.805 0.115 / 0.106 0.761 / 0.806
RSTN (Wang et al., 2023) 7.963 / 7.533 21.034 / 22.065 0.111 / 0.107 0.763 / 0.813
LGUTransformer (Wang et al., 2024b) 7.357 / 7.352 22.108 / 22.111 0.105 / 0.104 0.812 / 0.818
Table 7: Training using synthetic smoke and evaluation using real smoke on Cholec80 (non-overlap videos).
Method BRISQUE $\downarrow$ FADE $\downarrow$ NIQE $\downarrow$
AODNet (Li et al., 2017) 12.551 / 14.814 0.440 / 0.484 10.449 / 9.434
GCANet (Chen et al., 2019) 7.795 / 11.566 0.407 / 0.403 9.866 / 9.331
De-smokeGCN (Chen et al., 2020) 4.857 / 5.575 0.384 / 0.376 10.885 / 9.569
MSBDN (Dong et al., 2020) 11.838 / 12.176 0.421 / 0.411 10.299 / 9.026
SVPNet (Wang et al., 2024a) 12.812 / 12.134 0.416 / 0.402 10.579 / 9.387
CGAN-DC (Chen et al., 2019) 10.258 / 9.945 0.413 / 0.404 9.517 / 9.039
MARS-GAN (Hong et al., 2023) 13.999 / 11.680 0.416 / 0.413 13.336 / 13.884
SSIM-PAN (Sidorov et al., 2020) 10.539 / 11.093 0.411 / 0.403 9.811 / 11.999
RSTN (Wang et al., 2023) 13.868 / 12.652 0.420 / 0.414 12.245 / 16.910
LGUTransformer (Wang et al., 2024b) 9.126 / 11.236 0.417 / 0.416 13.734 / 12.471
Table 8: Training on VASST-desmoke (real smoke) and evaluation on Hamlyn Dataset (real smoke).
Method BRISQUE $\downarrow$ FADE $\downarrow$ NIQE $\downarrow$
AODNet (Li et al., 2017) 36.549 / 30.515 0.331 / 0.322 12.418 / 11.851
GCANet (Chen et al., 2019) 13.354 / 26.555 0.324 / 0.294 11.512 / 10.983
De-smokeGCN (Chen et al., 2020) 33.236 / 28.007 0.329 / 0.325 29.448 / 18.802
MSBDN (Dong et al., 2020) 26.992 / 25.198 0.384 / 0.322 14.224 / 10.944
SVPNet (Wang et al., 2024a) 39.862 / 30.325 0.401 / 0.317 16.245 / 13.306
CGAN-DC (Chen et al., 2019) 14.052 / 16.699 0.352 / 0.336 16.616 / 13.216
MARS-GAN (Hong et al., 2023) 33.722 / 30.601 0.377 / 0.369 15.011 / 16.381
SSIM-PAN (Sidorov et al., 2020) 11.936 / 16.047 0.357 / 0.298 16.613 / 20.297
RSTN (Wang et al., 2023) 35.629 / 30.232 0.412 / 0.330 16.442 / 12.645
LGUTransformer (Wang et al., 2024b) 35.260 / 29.706 0.295 / 0.294 15.300 / 12.135
Table 9: Training on Cholec80 (synthetic smoke) and evaluation on Hamlyn Dataset (real smoke).
Method BRISQUE $\downarrow$ FADE $\downarrow$ NIQE $\downarrow$
AODNet (Li et al., 2017) 30.292 / 31.168 0.426 / 0.432 11.074 / 11.834
GCANet (Chen et al., 2019) 19.366 / 26.102 0.385 / 0.293 11.628 / 11.083
De-smokeGCN (Chen et al., 2020) 13.920 / 17.647 0.330 / 0.315 11.628 / 11.467
MSBDN (Dong et al., 2020) 26.630 / 27.559 0.409 / 0.332 12.152 / 11.882
SVPNet (Wang et al., 2024a) 29.065 / 28.375 0.406 / 0.267 14.358 / 12.292
CGAN-DC (Chen et al., 2019) 25.175 / 24.127 0.408 / 0.315 10.445 / 10.741
MARS-GAN (Hong et al., 2023) 29.811 / 27.287 0.409 / 0.311 16.443 / 15.136
SSIM-PAN (Sidorov et al., 2020) 24.605 / 26.001 0.395 / 0.312 10.852 / 19.337
RSTN (Wang et al., 2023) 26.305 / 28.347 0.414 / 0.328 13.875 / 14.217
LGUTransformer (Wang et al., 2024b) 23.974 / 26.497 0.383 / 0.346 16.786 / 17.448
4.3 Comparison of Quantitative Results
We train five folds of ten existing methods and their corresponding SurgiATM modified versions, using Cholec80 with synthetic smoke and VASST-desmoke with real-world smoke, respectively. We then compare their five-fold average performance on the real-world surgical smoke datasets: VASST-desmoke, Cholec80 (real smoke), and the Hamlyn Dataset. In total, we train 200 models and conduct 500 evaluations.
The quantitative results on VASST-desmoke and VASST-desmoke-C are presented in Tables 4, 5 and 6. They demonstrate that incorporating SurgiATM effectively improves model performance, regardless of the backbone architecture or loss function. However, for SVPNet in Table 6, SurgiATM yields no improvement in SSIM. We argue that, compared to other desmoking methods, SVPNet (Wang et al., 2024a) employs an energy function (Wang et al., 2018) for restoration. This is a physics-guided approach similar to ours, and its three physics-related parameters are predicted by three independent U-Nets. As a result, incorporating an extra physics-guided component (our SurgiATM) into SVPNet may slightly disrupt the internal physical consistency, leading to the observed deterioration in SSIM.
Figure 5: Each heatmap illustrates the ablation study for a specific baseline model incorporated with SurgiATM, showing the impact of different hyperparameter settings in terms of the RMSE metric (lower is better). Each heatmap uses an independent color scale.
For Cholec80, we report the evaluation results in Table 7. Our SurgiATM effectively improves most model performances in terms of the fog-related metric (i.e., FADE score). In terms of the external evaluation on the Hamlyn Dataset, as shown in Tables 9 and 8, for the defogging metric FADE, employing our SurgiATM enhances most methods, except for the traditional dehazing method AODNet in Table 9. We attribute this to two factors: 1) the simple architecture and insufficient trainable weights of AODNet (i.e., only five convolutional layers) compared to other methods, and 2) the non-Laplacian error distribution of AODNet when trained on Cholec80 with synthetic smoke, as reported in Table 2. In contrast, although GCANet and SSIM-PAN also exhibit non-Laplacian error distributions in Cholec80, and CGAN-DC and MARS-GAN in VASST-desmoke, their performances are effectively enhanced owing to both their well-designed architectures and the strong generalizability of SurgiATM.
On the other hand, SurgiATM demonstrates limited effectiveness in terms of BRISQUE and NIQE across the ten existing methods in Tables 7, 8 and 9. The main reason is that these two metrics are designed to measure similarity to natural images, which may not generalize well to surgical endoscopic frames due to the substantial domain gap between them (Wang et al., 2024b). The scarcity of domain relevent non-referenced IQA metrics is exactly an underexplored research problem in the area of surgical smoke removal. However, addressing this limitation is beyond the scope of the current study, and developing reliable, domain-specific metrics for surgical videos remains a promising direction for future research.
4.4 Comparison of Qualitative Results
In this section, we present qualitative comparisons across several representative baselines: De-smokeGCN, GCANet, MARS-GAN, and SVPNet, each selected for different reasons. Specifically, De-smokeGCN (Chen et al., 2020) is a classic DNN-based desmoking method introduced in 2020; GCANet (Chen et al., 2019) is designed for natural dehazing and deraining; MARS-GAN represents an adversarial-learning-based approach; and SVPNet (Wang et al., 2024a), published in 2024, uses a physics-based loss function tailored for surgical desmoking. We train the models on Cholec80 with synthetic smoke and evaluate them on VASST-desmoke, Cholec80, and the Hamlyn Dataset with real surgical smoke, challenging their stability and generalizability.
As shown in the error maps in Fig. 7, SurgiATM effectively reduces desmoking errors in a computationally efficient and interpretable manner, making it well-suited for deployment on edge devices and for stable surgical desmoking. For the natural dehazing method GCANet, it causes color distortion in Fig. 7(a), which is corrected to a certain degree by using SurgiATM. In addition, the proposed approach can mitigate the grid artifact (see the first two rows in Fig. 7), demonstrating promising stability. As shown in Fig. 8, the desmoking results of the baselines are improved to varying degrees. For instance, De-smokeGCN and GCANet can produce more stable results with SurgiATM by reducing the grid artifacts (see Fig. 8(a)-(d) of De-smokeGCN) and correcting the restoration colors (see Fig. 8(d) of GCANet). Moreover, SurgiATM helps baselines eliminate more smoke, as shown in Fig. 8(a) for MARS-GAN and Fig. 8(c) for SVPNet, GCANet, and MARS-GAN.
5 Discussion
5.1 Ablation Study
We conduct ablation studies to investigate the influence of two hyperparameters (i.e., smoothing factor $\eta$ and shifting window size $z$) on model performance. Experiments are carried out on the real-world benchmark VASST-desmoke using three representative baselines (i.e., AODNet, RSTN, and SSIM-PAN), chosen for their minimal reliance on auxiliary modules and loss functions, thereby enabling a clear assessment of the impact of the only two hyperparameters, smoother $\eta$ and window size $z$, with minimal confounding factors. The RMSE metrics are averaged from five-fold cross-validation and reported in Fig. 5.
We observed that, for most methods, $\eta=0.1$ usually leads to better performance, a trend that is obviously observed for AODNet in Fig. 5. This hyperparameter defines a lower bound of denormalized dark channel $\mathcal{D}$ (see Eq. (29)) and consequently, sets the minimum gradient propagated from loss functions (see Eqs. (27) and (28)). Values that are too small values (e.g., 0.05, 0.01, or 0.001) can contribute to gradient vanishing, which hinders models fitting and updating. Conversely, large value allow higher-magnitude gradients, but the effectiveness of $\mathcal{D}$ is diminished (see Eq. (29)), gradually degenerating into general residual learning.
Another potential trend is observed with respect to the window size $z$: larger window sizes are preferred for AODNet and SSIM-PAN, whereas smaller ones perform better for RSTN. Analyzing their architectures, we infer that CNN backbones used by AODNet and SSIM-PAN, which have comparatively smaller receptive field than SwinT backbone used by RSTN, benefit from larger window size to capture visual information over a wider range.
Therefore, when using SurgiATM, $\eta=0.1$ is recommended as a default configuration, while the optimal window size should be selected based on the specific baseline method.
Figure 6: The average improvement of performance by using SurgiATM in terms of RMSE metrics facing different smoke density. The data is analyzed pixel-by-pixel from VASST-desmoke dataset, because it contains ground truth and supports pixel-wise measurement. ”Smoke Density” is exactly approximated with dark channel. The smoke images are generated by Perlin noise for illustrating the smoke density. ”Trimmed Mean” computes the mean improvement excluding the max and min values.
5.2 Performance Improvement Analysis
To further investigate the effectiveness of SurgiATM under varying smoke density, we analyze the performance improvements across all baseline methods and show the average enhancement trend in Fig. 6.
Overall, SurgiATM enables the baseline methods to generate more accurate frames to varying degrees. Specifically, it provides substantial improvements under thin-smoke conditions and moderate enhancements under dense smoke conditions. An inflection point is observed at a smoke density of approximately 0.5, beyond which the average error reduction stabilizes at around 7.5%. The performance drop observed at a smoke density of 1.0 is expected and can be attributed to the fact that, in such heavily smoke-contaminated regions, the dark channel value also approaches 1.0. Under this condition, SurgiATM degenerates into ordinary residual learning, making observed decline theoretically inevitable. The significance improvements under thin-smoke is mainly benifited from the ability of consistent predictoin that involves minimal or even no color shifting when $\mathcal{D}$ is very small or close to zero, which is more stable and trustworthy compared to learning a pure non-linear mapping with neural networks; this is also a main motivation for using Eq. (20).
Table 10: Average Incremental Overhead (%) by Using SurgiATM.
Method $\Delta$ Runtime $\Delta$ FLOPs
AODNet (Li et al., 2017) 2.0759 0.5244
CGAN-DC (Salazar-Colores et al., 2020) 0.6908 0.0043
De-smokeGCN (Chen et al., 2020) 0.2634 0.0057
GCANet (Chen et al., 2019) 0.3382 0.0038
LGUTransformer (Wang et al., 2024b) 0.7762 0.0011
MARS-GAN (Hong et al., 2023) 0.4748 0.0015
MSBDN (Dong et al., 2020) 0.5195 0.0025
RSTN (Wang et al., 2023) 0.4250 0.0050
SSIM-PAN (Sidorov et al., 2020) 0.6661 0.0102
SVPNet (Wang et al., 2024a) 0.5042 0.0011
Avg. 0.6734 0.0560
5.3 Computation Overhead Analysis
SurgiATM is a lightweight module, featured with only two hyperparameters and no additional trainable weights. Consequently, its computatinoal cost is minimal, requiring approximately $35$ microseconds ($\mu s$) and $9.8\times 10^{5}$ floating-point operations (FLOPs) per frame on our device setup (1 RTX 3090 GPU) with an input resolution of $256\times 256$. We compute the average runtime across all experiments, and report the incremental overhead introduced by SurgiATM in Table 10. These results demonstrate that SurgiATM has negligible impact on the inference speed of the integrated methods, with overall efficiency primarily determined by the baseline models themselves. This efficiency stems from the fact that SurgiATM is derived from classic physical models and statistical analysis without involving learnable components such as Multilayer Perceptrons (MLPs), thereby introducing minimal parameter overhead.
5.4 Downstream Task Analysis
High-fidelity surgical scene is a fundamental prerequisite for clinial applications that are largely affected by image quality (Zhao et al., 2025), and smoke removal is expected to support intelligent assistance functions such as surgical scene understanding. To assess the impact of SurgiATM on downstream tasks, we adopt MedSAM (Ma et al., 2024b), a foundation model for medical scene segmentation, to detect and delineate surgical instruments, under different smoke density and light conditions.
As shown in Fig. 9, all SOTA methods yield cleaner images and enhance the segmentation performance compared to directly using MedSAM. Moreover, by using SurgiATM, their results are further stabilized and refined to varying degrees, leading to more accurate segmentation on restored images. The results suggesting that SurgiATM supports other SOTA desmoking methods to deliver the higher overall segmentation accuracy, demonstrating its effectiveness in retaining task-relevant semantic cues that are essential for downstream tasks.
5.5 Integration Conflict Analysis
In Table 6, for SVPNet (Wang et al., 2024a), a physics-guided method, SurgiATM causes a slight drop for SSIM score, and minor enhancement on other scores. We investigate the primary mechanism of SVPNet and locate a problem. SVPNet used smoke veil $F\in\mathbb{R}^{W\times H}$ to model the observed frame:
$\ I(x,c)=F(x)+t(x,c)\cdot J(x,c).$ (30)
Different from other desmoking methods, SVPNet employed three seperated U-Nets to estimate $F$, $t$, and $J$ (i.e., multiple outpus), which conflicts with SurgiATM (i.e., single output). When SurgiATM is forcibly integrated with SVPNet (i.e., combining Eqs. (30) and (23)), the resulting formula becomes unintuitive and difficult to interpret:
$\ I(x,c)=\frac{F(x)-t(x,c)\cdot\mathcal{D}(x)\cdot(1-\rho(x,c))}{1-t(x,c)}.$ (31)
Although, Eq. (31) is differentiable with respect to $F$, $t$, and $\rho$, allowing the neural networks to be updated via backpropagation, the model outputs deviated from its original physical meaning. As a result, it becomes unclear what the networks are actually learning. We hypothesize that this loss of physical interpretability is a main reason why SurgiATM can produce unexpected effects when applied to SVPNet. Consequently, when using SurgiATM, baseline methods with single output are preferred. Forcibly merging SurgiATM with multi-output approaches may lead to unpredictable or undesired impact (i.e., insignificant improvement or slight decline).
6 Conclusion and Future Work
In this work, we propose SurgiATM for laparoscopic surgical smoke removal, which can be seamlessly incorporated into existing DNN-based desmoking methods without additional trainable parameters. By using SurgiATM, the performance and generalizability of these methods are enhanced, enabling the generation of cleaner frames for both surgeons and downstream tasks. We analyze the discrepancy between natural dehazing and surgical desmoking and design SurgiATM by first formulating a Mixture-of-Experts model. We then conduct a comprehensive statistical analysis of existing desmoking methods and derive a new restoration formula tailored for surgical smoke removal. Extensive experiments across three distinct datasets validate the effectiveness of SurgiATM.
Nevertheless, several limitations accompany its advantages. First, the indicator $W$ is currently estimated as a linear function of the dark channel $D$, which captures the general trends across methods, and thus leaves some room for improvement when focusing on a particular method; this could potentially be addressed by deriving method-specific formulations. Second, the denormalized dark channel obtained from DCP remains relatively coarse-grained. Developing an appropriate refinement strategy may further improve its performance and result quality; for example, using guided image filtering or attention mechanisms could generate a comparatively fine-grained denormalized dark channel.
Appendix A Visualization Results
Figure 7: Error comparison in VASST-desmoke (Xia et al., 2024). “+” indicates the baseline integrated with our SurgiATM. ”GT” represents ground truth. The methods are trained in Cholec80 (Twinanda et al., 2017) with synthetic smoke. Orange boxes mark the errors declined by SurgiATM. Overall, “Baseline + SurgiATM” exhibits more accurate restoration. ”a” shows a frame with addequate light covered by partial dense smoke; ”b” displays relatively thin smoke with severe light reflection; ”c” represents a frame with both dense smoke and light reflection. The error map demonstrates that SurgiATM can enhance the desmoking results facing to different conditions.
Figure 8: Comparison of SurgiATM with baselines on real datasets: Cholec80 (”a” and ”b”) and the Hamlyn Dataset (”c” and ”d”). Blue boxes highlight the comparatively accurate and stable restoration achieved by SurgiATM.
Figure 9: Qualitative comparisons on VASST-desmoke dataset with downstream segmentation task. The top row shows input frames and their segmentation before desmoking. ”Baseline” and ”Baseline+” represent baseline methods and SurgiATM’s version. All segmentation masks are generated by the foundation model MedSAM (Ma et al., 2024b). (a) shows a comparatively simple scenario with thin smoke, clear instrument, and plain background; (b) demonstrates a complexer one where the edge of the instrument is slightly blurred with background; (c) displays a challenging scene with heavy smoke and severe light reflection, where the instrument at the left side is barely visible.